### v1

# Pólya’s conjecture up to $\epsilon$-loss and quantitative estimates for the remainder of Weyl’s law 00footnotetext: 2020 Mathematics Subject Classification. Primary 35P15; Secondary 35P20, 42B37. Key words and phrases: Dirichlet eigenvalue, Weyl law, Pólya conjecture

Abstract. Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain. For any $\epsilon\in(0,1)$ we show that for any Dirichlet eigenvalue $\lambda_{k}(\Omega)>\Lambda(\epsilon,\Omega)$, it holds
$k$ $\leq(1+\epsilon)\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},$
where $\Lambda(\epsilon,\Omega)$ is given explicitly. This reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem. This estimate is based on a uniform estimate on the remainder of the Weyl law on Lipschitz domains, which appears to be the first quantitative estimate for the remainder of Weyl’s law since Weyl’s seminal work in the year 1911. We also provide in all dimensions $n\geq 2$, a class of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
1 Introduction
1.1 Weyl’s law and Pólya’s conjecture
Given an open bounded domain $\Omega$ in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$. Let
$0<\lambda_{1}(\Omega)\leq\lambda_{2}(\Omega)\leq\lambda_{3}(\Omega)\leq\cdots$
be eigenvalues to the Dirichlet Laplacian on $\Omega$. We shall simply write $\lambda_{k}$ instead of $\lambda_{k}(\Omega)$ when there is no confusion.
In 1911, Weyl [34] proved the famous Weyl’s law, which states that for any bounded open domain $\Omega\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$, and $\lambda\to\infty$,
(1.1) $\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+o(\lambda^{n/2}),$
where $\omega(n)$ denotes the volume of unit ball in ${{{\mathbb{R}}}^{n}}$, $\mathcal{N}^{D}_{\Omega}(\lambda)$ is the counting function
$\mathcal{N}^{D}_{\Omega}(\lambda):=\#\{\lambda_{k}(\Omega):\,\lambda_{k}(\Omega)<\lambda\}.$
The Weyl law is also important in the longstanding question “can one hear about the shape of a cavity”, see Lord Rayleigh [27] and Kac [21]. Moreover, the Weyl law is even related to the number theory, Prime number theorem, Riemann hypothesis, via Connes’ trace formula, see Connes [4], Fathizadeh-Khalkhali [11] and the beamer by Khalkhali [22].
Let us denote the remainder of Weyl’s law as
(1.2) $R_{\Omega}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$
Subsequent to Weyl’s seminal work, Courant [5], Courant-Hilbert [6], Hörmander [17, 18], Seeley [32, 33] and many others (see Ivrii [20] for comprehensive description on related history) made contributions to improve the estimate on the remainder. Up to present, the best estimate is as $O(\lambda^{\frac{n-1}{2}})$ on piecewise smooth domains by Seeley [32, 33], see also Hörmander [18].
Weyl also conjectured a sharper asymptotic behavior of Dirichlet eigenvalue states for domains with piecewise smooth boundary that
(1.3) $\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}-\frac{|\partial\Omega|}{2^{n+1}\pi^{\frac{d-1}{2}}\Gamma(\frac{n+1}{2})}\lambda^{\frac{n-1}{2}}+o(\lambda^{\frac{n-1}{2}}),$
here and in what follows, for any set $\Omega$, $|\partial\Omega|$ denotes the surface area of it, $\mathscr{B}(\cdot,\cdot)$ the Beta function and $\Gamma(\cdot)$ the Gamma function, as usual. The conjecture has been proved by Ivrii [19] under the condition that the set of all periodic geodesic billiards in $\Omega$ has measure zero, which was proved by Safarov-Vassiliev [31] for convex analytic domains and polygons.
However, none of the above mentioned results gave quantitative estimate for the remainder $R_{\Omega}(\lambda)$ (that is, these results states that certain asymptotic estimates valid for large $\lambda$’s without quantified the sizes of such $\lambda$’s), and consequently they do not give any quantitative estimate on eigenvalues. From both mathematical and physical point of view, it is rather natural to ask the following question.
Question 1.1.
Is it possible to give a uniform and quantitative estimate for the remainder $R_{\Omega}(\lambda)$ of Weyl’s law?
From the mathematical point of view, uniform and quantitative estimates are always more interesting. In fact, in 1954, Pólya [29] conjectured that, it should hold on arbitrary bounded open domain that
$k\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}^{n/2},\ \forall\,k\in{\mathbb{N}}.$
Pólya himself [30] proved this conjecture for tiling domains in the plane. His method extends to high dimensions, and so the conjecture is true for tiling domains in high dimensions too.
The Pólya conjecture has been wide open since then. Other than tiling domains, Laptev [24] verified Pólya’s conjecture on product domains $\Omega_{1}\times\Omega_{2}\subset{\mathbb{R}}^{n_{1}+n_{2}}$, $n_{1}\geq 2$, $n_{2}\geq 1$, provided Pólya’s conjecture holds on $\Omega_{1}$. Filonov et al. [7] and [8] verified, via a method from number theory, Pólya’s conjecture on balls for all $n\geq 2$, and on annuli for $n=2$. For general domains $\Omega$, it was known that Pólya’s conjecture holds for the first two Dirichlet eigenvalues see [16] for the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$). See also [16] for the first nontrivial Neumann eigenvalue and recent exciting developments [3, 13] for the second nontrivial Neumann eigenvalue.
For general domains, Berezin [2] and Li-Yau [26] independently proved for arbitrary bounded open domains that
(1.4) $\frac{|\Omega|^{2/n}\omega(n)^{2/n}}{(2\pi)^{2}}\sum_{i=1}^{k}\lambda_{i}\geq\frac{n}{n+2}k^{\frac{n+2}{n}},$
which implies that
(1.5) $\left(\frac{n+2}{n}\right)^{n/2}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}^{n/2}\geq k,$
see also Laptev [24], and [9, 10, 12, 23, 28] for further improvements on the lower order term and generalizations. Note that the estimate (1.5) and its improvements are still a distance away from Pólya’s conjecture.
Our main purpose of this paper is to prove Pólya’s conjecture up to an $\epsilon$-loss, and to give a solution to Question 1.1, on Lipschitz domains.
1.2 Main results
Let us first give a definition for Pólya’s conjecture up to an $\epsilon$-loss.
Definition 1.2.
Let $\Omega\subset{{{\mathbb{R}}}^{n}}$ be a bounded open domain, $n\geq 2$. For $\epsilon\in(0,1)$, we say that Pólya’s conjecture up to an $\epsilon$-loss or $\epsilon$-loss version of Pólya’s conjecture holds for some or all Dirichlet eigenvalues $\lambda_{k}(\Omega)$, if it holds for some or all $k\in{\mathbb{N}}$ that
$k$ $\leq(1+\epsilon)\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
Noting that $2\leq\left(\frac{n+2}{n}\right)^{n/2}\to e$ as $2\leq n\to\infty$, the $\epsilon$-loss version of Pólya’s conjecture improves substantially (1.5) and its improvements on lower order terms, see [2, 12, 23, 28].
For a bounded open domain $\Omega$, we define its width $\mathrm{width}(\Omega)$ as the minimum distance between two parallel hyperplanes that contain $\Omega$ inside. Moreover, up to a rotation, throughout the paper we shall assume the width $\mathrm{width}(\Omega)$ is attained in the $n$-th direction. For a bounded Lipschitz domain $\Omega$, we let $C_{Lip}(\Omega)\geq 1$ be the constant, uniquely determined by $\Omega$, such that for $\epsilon\in(0,\mathrm{width}(\Omega))$, such that
(1.6) |$\left|\{x\in{{{\mathbb{R}}}^{n}}:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)\leq\epsilon\}\right|\leq C_{Lip}(\Omega)\epsilon|\partial\Omega|.$
Our first main result is as following.
Theorem 1.3.
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded Lipschitz domain, $n\geq 2$. For any $10<\epsilon<$ 1, it holds for all $\lambda_{k}(\Omega)\geq\Lambda(\epsilon,\Omega)$ that
$k$ $\leq(1+\epsilon)\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},$
where $\Lambda(\epsilon,\Omega)\geq\mathrm{width}(\Omega)^{-2}$ is such that
(1.7) $\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}C_{Lip}(\Omega)\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
Obviously, for a given $\epsilon\in(0,1)$, and a concrete Lipschitz domain $\Omega$, $\Lambda(\epsilon,\Omega)$ is uniquely determined by the equation (1.7). Hence, to show the $\epsilon$-loss version of Pólya’s conjecture, Theorem 1.3 reduces the problem to a computational problem for eigenvalues $\lambda_{k}(\Omega)<\Lambda(\epsilon,\Omega)$.
Theorem 1.3 will follow from our following uniform estimate for the remainder of the Weyl’s law. Let us start from the following definition.
Definition 1.4 (Minimal Admissible rectangle).
Let $\mathscr{R}=I_{1}\times I_{2}\times\cdots\times I_{n}$ be a rectangle in ${\mathbb{R}}^{n}$, $n\geq 2$, where $I_{i}$ is connected open interval in ${\mathbb{R}}$ for each $1\leq i\leq n$. We say that $\mathscr{R}=I_{1}\times I_{2}\times\cdots\times I_{n}=\mathscr{R}_{n-1}\times I_{min}$ is the minimal admissible rectangle for $\Omega$, if $I_{min}=I_{n}=\mathrm{width}(\Omega)$, and $\mathscr{R}$ contains $\Omega$, and for each $i=1,\cdots,n$, for any $\tilde{I}_{i}$ with $|\tilde{I}_{i}|<|I_{i}|$, $I_{1}\times\cdots\tilde{I}_{i}\times\cdots\times I_{n}$ does not contain $\Omega$ inside.
We have the following uniform estimate on Lipschitz domains, which gives a solution to Question 1.1. Moreover, our uniform estimate appears to be the first quantitative estimate for the remainder of Weyl’s law since Weyl’s seminal work [34] in the year 1911, see also Courant [5], Courant-Hilbert [6], Hörmander [18], Ivrii [19], Seeley [32, 33] etc for asymptotic results. In what follows, for a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$.
Theorem 1.5.
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded Lipschitz domain, $n\geq 2$. Let $\mathscr{R}=\mathscr{R}_{n-1}\times I_{min}$ be the minimal Admissible rectangle for $\Omega$. It holds for all $k\in{\mathbb{N}}$ that
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}+C(n,\Omega,\lambda_{k}(\Omega))\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},$
where
(1.8) $C(n,\Omega,\lambda_{k}(\Omega))=C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\overline{\Omega})|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|.$
Moreover, it holds for each $k\in{\mathbb{N}}$ that
$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
Above,
(1.9) $C_{1}(n)=\begin{cases}\frac{2}{3}&n=1,\\ \frac{3\pi}{16}&n=2,\\ \frac{\pi\mathscr{B}(\frac{n}{2}-1,2)}{6\mathscr{B}(\frac{n}{2}-1,5/2)}&n=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n}{2}-1,2)}{16\mathscr{B}(\frac{n}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
Remark 1.6.
If $\Omega\subset{{{\mathbb{R}}}^{n}}$ is a convex set, then one can take $C_{Lip}(\Omega)$ to be 1 in the above two theorems. In fact, one can have better estimates than (1.7) and (1.8). See Remark (4.2) below.
Obviously, both quantities $\frac{r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}$ and $\frac{r_{in}(\Omega)}{|I_{min}|}$ are not larger than $1/2$, and therefore can be replaced by $1/2$. Moreover, note that if $|\partial(\mathscr{R}\setminus\overline{\Omega})|$ is small compared to $|\mathscr{R}_{n-1}|$, we then have that Pólya’s conjecture holds for the first some eigenvalues. Our quantitative estimate is consistent to the asymptotic estimates established by Courant [5] and Courant-Hilbert [6], which is, up to present, best possible in the case of domains having only Lipschitz continuous boundary.
Let us remark that the logarithm term in (1.8) comes from the Whitney decomposition for $\mathscr{R}\setminus\overline{\Omega}$, see the proof of Theorem 1.5 in Section 4 below. For an explicitly given Lipschitz domain $\Omega$, one may perform a better decomposition of $\mathscr{R}\setminus\overline{\Omega}$, which is not necessarily of Whitney type, and which may give better constant in the lower order terms if one can decompose the given domain into fewer and larger cubes or rectangles. Generally, to get better estimates than (1.8), one has to find some generic and better decomposition than the Whitney decomposition we used in the proof. If $\mathscr{R}\setminus\overline{\Omega}$ admits a better decomposition, such as a union of up to infinite many cubes with finite surface, we then can prove full Pólya’s conjecture, see discussions in the next subsection.
1.3 Pólya’s conjecture on a class of domains
Using the strategy developed above, we can provide in all dimensions $n\geq 2$, a class of domains which may have rather irregular shape or boundary but satisfy Pólya’s conjecture.
In what follows, by a cube $Q$ we always mean it is an open cube and denote by $\ell(Q)$ its side length. Let us first introduce a definition. Let us emphasize that, since the Pólya conjecture is scaling invariant, it is not restrictive to require in the following definition that the side length of the largest cube $Q$ is one.
Definition 1.7 (Admissible class).
Let $n\geq 2$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be a sequence of cubes in ${{{\mathbb{R}}}^{n}}$ with $\sup_{Q_{k}}\{\ell(Q_{k})\}=1$. Moreover, we allow some of these cubes $Q_{k}$ to be empty set. Set $\mathscr{S}_{Q}$ as the surface area of these cubes, i.e.,
$\mathscr{S}_{Q}:=\sum_{k\in{\mathbb{N}}}\ell(Q_{k})^{n-1}.$
We say that $\{Q_{k}\}_{k\in{\mathbb{N}}}$ is of Admissible class if $\mathscr{S}_{Q}<\infty$, and for any $k,j\in{\mathbb{N}}$, $Q_{k}\cap Q_{j}=\emptyset$ if $k\neq j$.
By a rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$, we mean $\mathscr{R}_{n-1}\subset{\mathbb{R}}^{n-1}$ is a rectangle, and $I\subset{\mathbb{R}}$ is an open interval.
Theorem 1.8.
Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\mathscr{R}$. If $|\mathscr{R}_{n-1}|>C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
(1.10) $C_{2}(n)=\begin{cases}2\sqrt{2}\pi,&n=1,\\ 9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
For $M\geq 2$, by $\tilde{\Omega}$ $M$-tiles $\Omega$ we mean there are $M$ isometries of $\tilde{\Omega}$, $\tilde{\Omega}_{1\leq k\leq M}$, with non-overlapped interior, such that $\cup_{1\leq k\leq M}\tilde{\Omega}_{k}\subset\Omega$ and $|\Omega|=M|\tilde{\Omega}|$.
Corollary 1.9.
Let $M\geq 2$. Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$, and $\Omega$ $M$-tiles $\mathscr{R}$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\Omega$. If $|\mathscr{R}_{n-1}|\geq C_{2}(n-1)M\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$.
Note that $M$ isometries of the domain $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$ corresponds to the case that removing $M$ copies of $\{Q_{k}\}_{k\in{\mathbb{N}}}$ from $\mathscr{R}$, the conclusion is obvious.
Some further remarks are in order. In what follows for $C>0$ we use $\lfloor C\rfloor$ to denote the integer part of it.
Remark 1.10.
(i) Let us take $\ell(k)=2^{\lfloor k/2\rfloor}$ for $k\geq 2$, $\ell(1)$=1, and $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ be a sequence of cubes, where $\ell(Q_{k,j})=2^{1-k}$. Then $\mathscr{S}_{Q}\leq 2(2+\sqrt{2})$. So for a rectangle $\Omega_{1}\times\Omega_{2}$, $|\Omega_{2}|>1$ and $|\Omega_{1}|>8(1+\sqrt{2})\pi$, one can arbitrarily remove any domain, that is composed by all or some of these cubes $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ with non-overlapped interior, from $\Omega_{1}\times\Omega_{2}$ such that the remaining set satisfies Pólya’s conjecture. See Figure 1 for some examples.
(a) Rectangle with a diamond and a room-corridor removed
(b) Trapezoid with infinite cubes removed
Figure 1: Examples satisfying Pólya’s conjecture
(ii) Previously, in ${\mathbb{R}}^{2}$, Pólya’s conjecture was only known to be true on two types of domains, tiling domains ([30]), balls, sectors and annuli ([7, 8]). In higher dimensions $n\geq 3$, other than tiling domains, balls and sectors that tile balls, there are a class of product domains satisfying Pólya’s conjecture ([24]), which are products of open sets in ${\mathbb{R}}^{n-k}$ and domains satisfying Pólya’s conjecture in ${\mathbb{R}}^{k}$ (tiling domains, balls or sectors that tile balls), $k\geq 2$. See also [15] for a specific product in ${\mathbb{R}}^{3}$.
Our result (Theorem 1.9) provides in all dimensions $n\geq 2$, a class of domains that may have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
1.4 Our method and plan of the paper
The main strategy to prove Theorem 1.3, Theorem 1.5 and Theorem 1.8 is as following.
We first use the explicit formula of Riesz means of Dirichlet eigenvalues in ${\mathbb{R}}$ (cf. Lemma 2.1), following the inspiring observation by Laptev [24], to give a refined estimates on product domains $\Omega_{1}\times\Omega_{2}$ (see Theorem 2.3 below). This allows us to conclude via monotonicity of Dirichlet eigenvalues (cf. [16]) that for any open domains $\Omega$ contained in $\Omega_{1}\times\Omega_{2}$, the first $k_{0}$ Dirichlet eigenvalues satisfies Pólya’s conjecture (cf. Corollary 2.4).
We then establish a quantitative lower bound for the counting function (upper bound for the Dirichlet eigenvalue) on cubes (see Lemma 3.2 and Lemma 3.5 below), which together with our refinements on product domains and monotonicity of Dirichlet eigenvalues to conclude the proof Theorem 1.8. In fact, in higher dimensions, $n\geq 3$, the rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I$ can be replaced some more general product $\Omega\times I$, as soon as Pólya’s conjecture holds on $\Omega$, see Theorem 3.6 below.
Although only a part of discussions in Sections 2 and 3 will be sufficient for the proofs of main results, the extended discussions are of independent interest and may be used for other purposes. For Theorem 1.5 we shall apply the Whitney decomposition of a Lipschitz domain $\Omega$, and using the lower bound for the counting function on cubes (see Lemma 3.2 and Lemma 3.5 below) along with an optimization to deduce a lower bound for the counting function $\mathcal{N}_{\Omega}^{D}(\lambda)$ on $\Omega$. Finally, applying our refinements on product domains (see Theorem 2.3 below) and monotonicity of Dirichlet eigenvalues, together with the lower bound of $\mathcal{N}_{\mathscr{R}\setminus\overline{\Omega}}^{D}(\lambda)$ to conclude the proof. Theorem 1.3 then follows almost immediately from Theorem 1.5.
Finally, let us point it out, our approach depends heavily on the min-max principle (monotonicity of Dirichlet eigenvalues), which is close to the original one by Courant [5] and Courant-Hilbert [6]. However, the way we use the min-max principle is rather different from that in Courant [5] and Courant-Hilbert [6]. By our choose of minimal Admissible rectangles, we do not even have to take care of the cubes which intersect the boundary of $\Omega$. The latter is the main issue for Courant [5] and Courant-Hilbert [6].
The paper is organized as follows.
In Section 2, we study the refined estimates on product domains, and prove the main results Theorem 2.3 and Corollary 2.4.
In Section 3, we provide the proof for the results on concrete examples of Pólya’s conjecture, Theorem 1.8. Since in higher dimensions, $n\geq 3$, we can replace the rectangle $\mathscr{R}_{n-1}$ by general set satisfying Pólya’s conjecture, we shall split the proof into the case of two dimension, and three and higher dimensions where we prove a more general result, Theorem 3.6. We shall also provide a qualitative result on removing general piecewise smooth set, Theorem 3.1, and shall show some results regarding cubes of Admissible class from some more general class of domains, see Theorem 3.7 below.
We shall employ Lemma 3.2 and Lemma 3.5 from Section 3 and Theorem 2.3 from Section 2 to prove the uniform estimates on Lipschitz sets, Theorem 1.5 and its corollary, Theorem 1.3. We shall also provide a uniform estimate on general open domains, see Theorem 4.3 below.
For $C>0$ we use $\lfloor C\rfloor$ to denote the integer part of it. For a cube $Q$ we denote by $\ell(Q)$ its side length. For a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$. Throughout the paper, most of constants are given explicitly, if not we shall point it out.
2 Refined eigenvalue estimates on product domains
In this section, we prove refined eigenvalue estimates on product domains, Theorem 2.3 and its corollary, Corollary 2.4. Recall that, if an open domain $\Omega$ is in the real line, then its eigenvalues have explicit lower bounds, which satisfy for $k\geq 1$
$\lambda_{k}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}}k^{2}.$
This allows us to show the following estimate for the Riesz means.
Lemma 2.1.
Let $\Omega\subset{\mathbb{R}}$ be bounded. It holds that
(2.1) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $\leq\frac{2|\Omega|}{3\pi}\lambda^{3/2}\min\left\{\left(1-\frac{3\pi}{16|\Omega|\sqrt{\lambda}}\right)_{+},\,\left(1-\left(\frac{\lfloor\frac{|\Omega|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega|\sqrt{\lambda}}\right)\right\},$
and for $\lambda>\frac{\pi^{2}}{|\Omega|^{2}}$,
(2.2) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\leq\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\leq\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
Moreover, if $\Omega$ is a connected open interval, it holds for $\lambda>\lambda_{1}(\Omega)=\frac{\pi^{2}}{|\Omega|^{2}}$ that
(2.3) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\geq\frac{|\Omega|}{4}\lambda\left(1-\frac{2\pi}{|\Omega|\sqrt{\lambda}}+\frac{\pi^{2}}{|\Omega|^{2}\lambda}\right).$
Proof.
Let us first prove (2.1). If $\lambda\leq\frac{\pi^{2}}{|\Omega|^{2}}$, then $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})=0$. Suppose $\lambda>\frac{\pi^{2}}{|\Omega|^{2}}$ and set $k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$. Then we have
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $=\sum_{1\leq k\leq k_{\lambda}}(\lambda-\lambda_{k})\leq\sum_{1\leq k\leq k_{\lambda}}(\lambda-\frac{\pi^{2}}{|\Omega|^{2}}k^{2})$
$=\lambda k_{\lambda}-\frac{\pi^{2}}{|\Omega|^{2}}\frac{1}{6}k_{\lambda}\left(k_{\lambda}+1\right)\left(2k_{\lambda}+1\right)$
$=\lambda k_{\lambda}-\frac{\pi^{2}}{3|\Omega|^{2}}k_{\lambda}^{3}-\frac{\pi^{2}}{2|\Omega|^{2}}k_{\lambda}^{2}-\frac{\pi^{2}}{6|\Omega|^{2}}k_{\lambda}.$
Since
$\lambda k_{\lambda}-\frac{\pi^{2}}{3|\Omega|^{2}}k_{\lambda}^{3}\leq\frac{2|\Omega|}{3\pi}\lambda^{3},$
we see that
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $<\frac{2|\Omega|}{3\pi}\lambda^{3}-\frac{\pi^{2}}{2|\Omega|^{2}}\left(\frac{k_{\lambda}}{\frac{|\Omega|}{\pi}\sqrt{\lambda}}\right)^{2}\frac{|\Omega|^{2}}{\pi^{2}}\lambda=\frac{2|\Omega|}{3\pi}\lambda^{3}-\left(\frac{k_{\lambda}}{\frac{|\Omega|}{\pi}\sqrt{\lambda}}\right)^{2}\frac{\lambda}{2}.$
For $\lambda>\frac{\pi^{2}}{|\Omega|^{2}}$, $\frac{k_{\lambda}}{\frac{|\Omega|}{\pi}\sqrt{\lambda}}\geq 1/2$, so we also have
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $<\frac{2|\Omega|}{3\pi}\lambda^{3}-\frac{\lambda}{8}.$
Let us prove (2.2). Again we only need to consider $\lambda>\frac{\pi^{2}}{|\Omega|^{2}}$. Set $k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$. Note that
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\leq\sum_{1\leq k\leq k_{\lambda}}(\lambda-\frac{\pi^{2}}{|\Omega|^{2}}k^{2})^{1/2}=\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}.$
Let us consider the ball $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. For each $1\leq k\leq k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$, $(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}$ equals the area of the rectangle $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, which is contained in $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. Moreover, for each $k$, there is at least a triangle with vertices at $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-(k-1)^{2})^{1/2})$, $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ and $(k,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ contained in the ball. Moreover, if $k_{\lambda}\neq\frac{|\Omega|}{\pi}\sqrt{\lambda}$, there is another triangle insider the 1/4 ball with vertices at $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$, $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$ and $(\frac{|\Omega|}{\pi}\sqrt{\lambda},0)$. See Figure 2111Picture drawn by Xinyi Jiang..
Figure 2: The 1/2 sum is controlled by the 1/4 ball
We therefore see that
$\frac{\pi}{4}\frac{|\Omega|^{2}}{\pi^{2}}\lambda$ $>\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}+\frac{1}{2}\sum_{1\leq k\leq k_{\lambda}}1\times\left((\frac{|\Omega|^{2}}{\pi^{2}}\lambda-(k-1)^{2})^{1/2}-(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}\right)$
$\quad+\frac{1}{2}\left(\frac{|\Omega|}{\pi}\sqrt{\lambda}-k_{\lambda}\right)\times(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2}$
$=\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}+\frac{1}{2}\left((\frac{|\Omega|^{2}}{\pi^{2}}\lambda)^{1/2}-(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2}+\left(\frac{|\Omega|}{\pi}\sqrt{\lambda}-k_{\lambda}\right)\times(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2}\right).$
For $x\in[0,1)$, let us consider the function
$g(x):=k_{\lambda}+x-\left((k_{\lambda}+x)^{2}-k_{\lambda}^{2}\right)^{1/2}(1-x).$
A calculation shows that for $k_{\lambda}\in{\mathbb{N}}$, it holds
$g(x)$ $=k_{\lambda}+x-\left(2k_{\lambda}x+x^{2}\right)^{1/2}(1-x)$
$\geq k_{\lambda}+x-\sqrt{2k_{\lambda}+1}\sqrt{x}(1-x)$
$\geq k_{\lambda}+x-\frac{2\sqrt{3}}{9}\sqrt{2k_{\lambda}+1}$
$\geq k_{\lambda}+x-\frac{2}{3}k_{\lambda}$
$\geq\frac{k_{\lambda}+x}{3}.$
This implies that
$\frac{\pi}{4}\frac{|\Omega|^{2}}{\pi^{2}}\lambda$ $>\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}+\frac{1}{2}\frac{|\Omega|}{\pi}\sqrt{\lambda}-\frac{\sqrt{6}}{9}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}$
$\geq\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}+\frac{1}{6}\frac{|\Omega|}{\pi}\sqrt{\lambda},$
i.e.,
$\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}<\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}=\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
Let us prove (2.3). Since $\Omega$ is connected, we have
$\lambda_{k}(\Omega)=\frac{k^{2}\pi^{2}}{|\Omega|^{2}},\,\forall\ k\in{\mathbb{N}}.$
Set again $k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$. For $\lambda>\frac{\pi^{2}}{|\Omega|^{2}}$, we have
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $=\sum_{1\leq k\leq k_{\lambda}}(\lambda-\frac{\pi^{2}}{|\Omega|^{2}}k^{2})^{1/2}=\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}.$
Note that the union of the rectangles $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, $1\leq k\leq k_{\lambda}$, contains the ball $B(0,\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda}+1})$ intersecting the first quadrant, since
$\sqrt{(k-1)^{2}+(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})}$ $=\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2k+1}\geq\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2k_{\lambda}+1}$
$\geq\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda}+1}.$
We therefore deduce that
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\geq\frac{\pi}{4}\frac{\pi}{|\Omega|}\left(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda}+1\right)$
$=\frac{|\Omega|}{4}\lambda\left(1-\frac{2\pi}{|\Omega|\sqrt{\lambda}}+\frac{\pi^{2}}{|\Omega|^{2}\lambda}\right).$
The proof is complete. ∎
Lemma 2.2.
Let $\Omega\subset{\mathbb{R}}$ be an open bounded set. Then for $p>1$ and $\lambda>\lambda_{1}=\lambda_{1}(\Omega)$, it holds that
$\sum_{k:\,\lambda_{k}<\lambda}(\lambda-\lambda_{k})^{p}$ $\leq\frac{2|\Omega|}{3\pi\mathscr{B}(p-1,2)}\lambda^{p+\frac{1}{2}}\mathscr{B}(p-1,5/2)-C_{3}(p)\lambda^{p},$
where
(2.4) $C_{3}(p)=\begin{cases}\frac{1}{8},&p=1,\\ \frac{1}{9},&1<p\leq 2,\\ \frac{1}{8}\left(\frac{247}{256}\right)^{p-1},&p>2.\end{cases}$
Proof.
By the semigroup type property of the Riesz means and (2.1), we have
[…]
Noticing that $\lambda_{1}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}},$ so for $\lambda>\lambda_{1}(\Omega)$,
$\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}\geq\frac{247}{256}\lambda.$
For $1<p\leq 2$,
[…]
For $p>2$, we have
[…]
$=\frac{\lambda^{p}}{p(p-1)}-\left(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}\right)^{p-1}\left(\frac{\lambda}{p-1}-\frac{(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}})}{p}\right)$
$\leq\frac{\lambda^{p}}{p(p-1)}-\left(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}\right)^{p-1}\frac{\lambda}{p(p-1)}$
and therefore
[…]
$\geq\frac{1}{8\mathscr{B}(p-1,2)}\left(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}\right)^{p-1}\frac{\lambda}{p(p-1)}$
$\geq\frac{\lambda^{p}}{8}\left(\frac{247}{256}\right)^{p-1}.$
Combining the above estimates yields the desired conclusion. ∎
In view of Laptev’s beautiful theory on product domains [24] (which has been further employed in Laptev-Weidl [25] for Schrödinger operator), the Aizenman-Lieb principle [1], we have the following refined eigenvalue estimates on product domains if one domain satisfies Pólya’s conjecture. Note that any bounded domain $\Omega\subset{\mathbb{R}}$ satisfies Pólya’s conjecture, i.e., $\lambda_{k}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}}k^{2}$ for each $k\in{\mathbb{N}}$.
Theorem 2.3.
Let $\Omega_{1}\subset{\mathbb{R}}^{n_{1}}$ and $\Omega_{2}\subset{\mathbb{R}}$ be bounded open domains, $n_{1}\geq 1$ and $n=n_{1}+1$. Suppose that Pólya’s conjecture holds on $\Omega_{1}$. Then it holds for $\lambda>0$ that
$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n_{1})}{|\Omega_{2}|\sqrt{\lambda}}\right)_{+},$
where $C_{1}(n_{1})$ is as in (1.9),
(2.7) $C_{1}(n_{1})=\begin{cases}\frac{2}{3}&n_{1}=1,\\ \frac{3\pi}{16}&n_{1}=2,\\ \frac{\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{6\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}&n_{1}=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{16\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n_{1}-2}{2}}&n_{1}\geq 5.\end{cases}$
Proof.
Suppose that it holds that
$\mathcal{N}^{D}_{\Omega_{1}}(\lambda)=\#\{\lambda_{k}(\Omega_{1}):\,\lambda_{k}(\Omega_{1})<\lambda\}\leq\frac{|\Omega_{1}|\omega(n_{1})}{(2\pi)^{n_{1}}}\lambda^{n/2}.$
Then for all $\lambda>\lambda_{1}(\Omega)$, where
(2.8) $\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{1})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{1}|^{2/n}\omega(n_{1})^{2/n_{1}}}+\frac{\pi^{2}}{|\Omega_{2}|^{2}},$
we have
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $=\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\#\left\{k:\,\lambda_{k}(\Omega_{1})+\lambda_{j}(\Omega_{2})<\lambda\right\}=\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\#\left\{k:\,\lambda_{k}(\Omega_{1})<(\lambda-\lambda_{j}(\Omega_{2}))_{+}\right\}$
$\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|\omega(n_{1})}{(2\pi)^{n_{1}}}(\lambda-\lambda_{j}(\Omega_{2}))^{n_{1}/2}.$
If $n_{1}=1$, then by (2.2), we have
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|}{\pi}(\lambda-\lambda_{j}(\Omega_{2}))^{1/2}\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{3|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
If $n_{1}=2$, then by (2.1) we have
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|\omega(2)}{(2\pi)^{2}}(\lambda-\lambda_{j}(\Omega_{2}))\leq\frac{|\Omega_{1}|\omega(2)}{(2\pi)^{2}}\frac{2|\Omega_{2}|}{3\pi}\lambda^{3/2}\left(1-\frac{3\pi}{16|\Omega_{2}|\sqrt{\lambda}}\right)_{+}$
$=\frac{|\Omega|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{3\pi}{16|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
For $n_{1}\geq 3$, by Lemma 2.2, we see that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|\omega(n_{1})}{(2\pi)^{n_{1}}}(\lambda-\lambda_{j}(\Omega_{2}))^{n_{1}/2}$
$\leq\frac{|\Omega_{1}|\omega(n_{1})}{(2\pi)^{n_{1}}}\left(\frac{2|\Omega_{2}|}{3\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}\lambda^{\frac{n_{1}}{2}+\frac{1}{2}}\mathscr{B}(\frac{n_{1}}{2}-1,5/2)-C_{3}(\frac{n_{1}}{2})\lambda^{\frac{n_{1}}{2}}\right)_{+}$
$=\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n_{1}+1}}\lambda^{\frac{n_{1}}{2}+1}\left(1-\frac{3\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{2\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}\frac{C_{3}(\frac{n_{1}}{2})}{|\Omega_{2}|\sqrt{\lambda}}\right)_{+}$
$=\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n_{1}+1}}\lambda^{\frac{n_{1}}{2}+1}\left(1-\frac{C_{1}(n_{1})}{|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
The proof is complete. ∎
As a corollary, we have the following result.
Corollary 2.4.
Let $\Omega_{1}\subset{\mathbb{R}}^{n_{1}}$ and $\Omega_{2}\subset{\mathbb{R}}$ be bounded open domains, $n_{1}\geq 1$ and $n=n_{1}+1$. Suppose that Pólya’s conjecture holds on $\Omega_{1}$. Let $\Omega$ be an open set contained in $\Omega_{1}\times\Omega_{2}$, and
$k_{0}=\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n_{1})}{|\Omega_{2}|}\right]^{n}\right\rfloor.$
Then for each $1\leq k\leq k_{0}$, the Pólya conjecture holds on $\Omega$ for $\lambda_{k}(\Omega)$, i.e.,
$k\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
Proof.
Since $\Omega\subset\Omega_{1}\times\Omega_{2}$, we may assume that $|\Omega|<|\Omega_{1}||\Omega_{2}|$, otherwise $|\Omega|=|\Omega_{1}||\Omega_{2}|$ and the Pólya conjecture holds for all Dirichlet eigenvalues of $\Omega$. Fix $\lambda_{0}>0$ such that
$|\Omega|=|\Omega_{1}||\Omega_{2}|\left(1-\frac{C_{1}(n_{1})}{|\Omega_{2}|\sqrt{\lambda}_{0}}\right).$
By the monotonicity of Dirichlet eigenvalues, for $\lambda<\lambda_{0}$ we have
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n_{1})}{|\Omega_{2}|\sqrt{\lambda}}\right)$
(2.9) $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$
Recall that
$k_{0}=\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n_{1})}{|\Omega_{2}|}\right]^{n}\right\rfloor\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{0}^{n/2}.$
Let $1\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
$k\leq\mathcal{N}^{D}_{\Omega}(\lambda_{k}(\Omega)+\epsilon)$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}(\lambda_{k}(\Omega)+\epsilon)^{n/2}.$
Letting $\epsilon\to 0$ gives that
$k\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
If $\lambda_{k}(\Omega)\geq\lambda_{0}$, then by the choose of $k_{0}$, we see that
$k\leq k_{0}\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{0}^{n/2}\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
The proof is complete. ∎
Let us consider $\Omega_{1}\subset{\mathbb{R}}^{2}$, with $|\Omega_{1}|=100$, $\Omega_{2}\subset{\mathbb{R}}$ with $|\Omega_{2}|=1$. Suppose that Pólya’s conjecture holds on $\Omega_{1}$ for the Dirichlet eigenvalues. In this case, $C_{1}(2)=3\pi/16$,
$\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n_{1})}{|\Omega_{2}|}\right]^{n}\geq\frac{|\Omega|9\pi}{2^{13}}\left(1-\frac{|\Omega|}{100}\right)^{-3}.$
For any $\Omega\subset\Omega_{1}\times\Omega_{2}$, if $|\Omega|\geq 95$, then
$\frac{|\Omega|9\pi}{2^{13}}\left(1-\frac{|\Omega|}{100}\right)^{-3}\geq 2623,$
and at least the first 2623 Dirichlet eigenvalues of $\Omega$ satisfy Pólya’s conjecture. If $|\Omega|\geq 99$ (resp. $|\Omega|\geq 80$), then at least the first 341694 (resp. 34) Dirichlet eigenvalues satisfy Pólya’s conjecture.
3 Pólya’s conjecture on a class of domains
Let us start with proving the qualitative result regarding removing general piecewise smooth set.
Theorem 3.1.
Suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture, $n\geq 2$, $\Omega_{2}$ is an interval in ${\mathbb{R}}$ and $\Omega_{3}$ is a piecewise smooth domain contained in $\Omega_{1}\times\Omega_{2}$. There exists $C_{P}=C(\Omega_{3},n)$ such that if $|\Omega_{1}|\geq C_{P}$ then the Pólya conjecture holds on $\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$.
Proof.
Let $\Omega=\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$. Up to a scaling, we may assume that $|\Omega_{3}|=1$. It follows from Seeley [32, 33] (see also [18, Corollary 17.5.11 & (17.5.16)’]) that for some $\Lambda>0$, it holds for $\lambda>\Lambda$ that
(3.1) $\left|\mathcal{N}^{D}_{\Omega_{3}}(\lambda)-\frac{|\Omega_{3}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|\leq C(\Omega_{3})\lambda^{\frac{n-1}{2}}.$
We assume that
$|\Omega_{1}|>\max\left\{\frac{C(\Omega_{3})(2\pi)^{n}}{\omega(n)C_{1}(n-1)},\frac{\sqrt{2\Lambda}}{C_{1}(n-1)}\right\}.$
By Corollary 2.4, Pólya’s conjecture holds for the first $k_{0}$ Dirichlet eigenvalues, where
$k_{0}=\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n-1)}{|\Omega_{2}|}\right]^{n}\right\rfloor=\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}(C_{1}(n-1)|\Omega_{1}|)^{n}\right\rfloor.$
It holds
$\lambda_{k_{0}}(\Omega)^{n/2}\geq\frac{(2\pi)^{n}}{|\Omega|\omega(n)}\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}(C_{1}(n-1)|\Omega_{1}|)^{n}\right\rfloor\geq\frac{1}{2}(C_{1}(n-1)|\Omega_{1}|)^{n},$
and hence,
$\lambda_{k_{0}}(\Omega)\geq\frac{1}{2}(C_{1}(n-1)|\Omega_{1}|)^{2}.$
Since $|\Omega_{1}|>\sqrt{2\Lambda}/C_{1}(n-1)$, it holds that
$\lambda_{k_{0}}(\Omega)>\Lambda.$
By (3.1) and $|\Omega_{3}|=1$, we see that for any $\lambda_{k}(\Omega)\geq\lambda_{k_{0}}(\Omega)$,
$\mathcal{N}^{D}_{\Omega_{3}}(\lambda_{k}(\Omega))\geq\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-C(\Omega_{3})\lambda_{k}(\Omega)^{\frac{n-1}{2}}.$
This implies on $\Omega_{3}$, there are $\mathcal{N}^{D}_{\Omega_{3}}(\lambda_{k}(\Omega))$ eigenvalues smaller than $\lambda_{k}(\Omega)$. Since $\Omega$ and $\Omega_{3}$ are disjoint, $\lambda_{k}(\Omega)$ corresponds to some $\lambda_{k+\tilde{k}}(\Omega\cup\Omega_{3})$ for some
$\tilde{k}\geq\mathcal{N}^{D}_{\Omega_{3}}(\lambda_{k}(\Omega))\geq\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-C(\Omega_{3})\lambda_{k}(\Omega)^{\frac{n-1}{2}}.$
Since $\Omega\cup\Omega_{3}\subset\Omega_{1}\times\Omega_{2}$ and $|\Omega\cup\Omega_{3}|=|\Omega_{1}\times\Omega_{2}|$, the monotonicity of Dirichlet eigenvalues together with Theorem 2.3 implies that
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|\Omega_{2}|\sqrt{\lambda}}\right)_{+},$
and therefore
$\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$ $=\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda_{k+\tilde{k}}(\Omega\cup\Omega_{3})^{n/2}$
$\geq k+\tilde{k}+\frac{C_{1}(n-1)}{|\Omega_{2}|\sqrt{\lambda}}\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}$
$\geq k+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-C(\Omega_{3})\lambda_{k}(\Omega)^{\frac{n-1}{2}}+\frac{C_{1}(n-1)|\Omega_{1}|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}.$
Since $|\Omega_{1}|\geq\frac{C(\Omega_{3})(2\pi)^{n}}{\omega(n)C_{1}(n-1)}$, we finally see that
$\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$ $=\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
$\geq k+\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{C_{1}(n-1)|\Omega_{1}|\omega(n)}{(2\pi)^{n}}-C(\Omega_{3})\right)$
$\geq k.$
The proof is complete. ∎
We next prove the concrete examples of domains obtained by removing cubes, Theorem 1.8. Let us start with the two dimensional case.
3.1 The case of two dimension
Let $Q\subset{\mathbb{R}}^{n}$ be an open unit cube, $n=2$. Then the Dirichlet eigenvalues of $Q$ has the form
$\left\{\lambda_{k}:\,\lambda_{k}=\pi^{2}(a^{2}+b^{2}),\,a,b\in\mathbb{N}\right\}.$
Note that for a given $\lambda_{k}$, $\mathcal{N}_{Q}^{D}(\lambda_{k})$ is dominated by the number of unit cubes contained in the first quadrant intersecting the ball $B(0,\sqrt{\lambda_{k}/\pi})$.
So it is readily to see that for all $k\in{\mathbb{N}}$ that $\lambda_{k}\geq 4\pi k.$ In fact Theorem 2.3 gives a better result. We however shall need the opposite estimates.
Lemma 3.2.
Let $Q\subset{\mathbb{R}}^{2}$ be a cube with side length $\ell(Q)$ in ${\mathbb{R}}^{2}$.
(i) For any $\lambda>2\pi^{2}/\ell(Q)^{2}$, it holds that
$\mathcal{N}_{Q}^{D}(\lambda)\geq\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\omega(2)2^{3/2}\pi\ell(Q)}{(2\pi)^{2}}\sqrt{\lambda}+\frac{\pi}{2}.$
(ii) For any $\lambda>0$, it holds that
$\mathcal{N}_{Q}^{D}(\lambda)\geq\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\omega(2)2^{3/2}\pi\ell(Q)}{(2\pi)^{2}}\sqrt{\lambda}.$
Proof.
We have
$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}),\,a,b\in\mathbb{N}\right\}.$
So the number $\mathcal{N}_{Q}^{D}(\lambda)$ is not less than the number of the unit cubes contained in the first quadrant intersecting the ball $B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{2})$, where $\sqrt{2}$ is the diameter of the unit cube. This implies that for any $\lambda>2\pi^{2}/\ell(Q)^{2}$,
$\mathcal{N}_{Q}^{D}(\lambda)$ $\geq\frac{1}{4}|B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{2})|=\frac{\pi}{4}\left(\ell(Q)\sqrt{\lambda}/\pi-\sqrt{2}\right)^{2}$
$\geq\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q)}{2}\sqrt{\lambda}+\frac{\pi}{2}.$
Note that for $\lambda\leq 2\pi^{2}/\ell(Q)^{2}$,
$\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q)}{2}\sqrt{\lambda}<0<\mathcal{N}_{Q}^{D}(\lambda),$
the conclusion is obvious. The proof is complete. ∎
Remark 3.3.
Using (2.3) one can give a lower bound for $\mathcal{N}^{D}_{\mathcal{R}}(\lambda)$ for any rectangle $\mathcal{R}$ in ${\mathbb{R}}^{2}$. Inductively, a lower bound for the eigenvalue counting function can be explicitly given on any cuboid in higher dimension.
Let us prove the planer case from Theorem 1.8.
Proof of the case $n=2$ in Theorem 1.8.
Denote by $\Omega_{0}=\Omega_{1}\times\Omega_{2}$ and $\Omega=\Omega_{1}\times\Omega_{2}\setminus\overline{\cup_{k=1}^{\infty}Q_{k}}$. Let
$\mathscr{V}_{Q}:=\left|\cup_{k=1}^{\infty}Q_{k}\right|=\sum_{k=1}^{\infty}\ell(Q_{k})^{n}.$
Then
$\mathscr{V}_{Q}\leq\mathscr{S}_{Q}<\infty.$
By Theorem 2.3 we have for $\lambda>0$ that
$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{{2}}{3|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
by Corollary 2.4 we have for
$k_{0}=\left\lfloor\frac{|\Omega|}{4\pi}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{2}{3|\Omega_{2}|}\right]^{2}\right\rfloor=\left\lfloor\frac{|\Omega|}{4\pi}\left[\frac{2|\Omega_{1}|}{3\mathscr{V}_{Q}}\right]^{2}\right\rfloor,$
the first $k_{0}$ Dirichlet eigenvalues of $\Omega$ satisfy Pólya’s conjecture, that is, for $1\leq i\leq k_{0}$,
$\frac{|\Omega|\omega(2)}{(2\pi)^{2}}\lambda_{i}(\Omega)\geq i.$
In particular, since $|\Omega_{1}|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}\geq 2\sqrt{2}\pi\mathscr{V}_{Q}$,
$|\Omega|=|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1)\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1),$
we find that
$\lambda_{k_{0}}(\Omega)\geq\frac{4\pi}{|\Omega|}\left(\frac{|\Omega|}{4\pi}\left[\frac{2|\Omega_{1}|}{3\mathscr{V}_{Q}}\right]^{2}-1\right)>33.$
Recall that Lemma 2.1 implies that for $\lambda>\pi^{2}/|\Omega_{1}|^{2}+\pi^{2}/|\Omega_{2}|^{2}$,
$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|}{\pi}(\lambda-\lambda_{j}(\Omega_{2}))^{1/2}$
$\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\lfloor\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right).$
In particular, for $\lambda>33$, it holds that
$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\frac{3}{2}\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}}\right)$
(3.2) $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{1}{|\Omega_{2}|\sqrt{\lambda}}\right).$
Now consider the union of open sets $\Omega$, $Q_{k}$, $k\in{\mathbb{N}}$, which are disjoint and contained in $\Omega_{1}\times\Omega_{2}$, but satisfy
$|\Omega_{1}\times\Omega_{2}|=|\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}|.$
So by monotonicity of Dirichlet eigenvalues, Pólya’s conjecture holds on $\Omega\cup\cup_{1\leq k<\infty}Q_{k}$,
$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{0}}(\lambda).$
Moreover, by Lemma 3.2 [(ii)], we see that for $\lambda\geq\lambda_{k_{0}}(\Omega)$,
$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}_{Q_{k}}(\lambda)$
$\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right).$
Combining the above two estimates, we deduce that for $\lambda\geq\lambda_{k_{0}}(\Omega)>33$,
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{{1}}{|\Omega_{2}|\sqrt{\lambda}}\right)-\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right)$
$=\frac{|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|\Omega_{1}|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right)$
$=\frac{|\Omega|}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|\Omega_{1}|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right).$
Since $|\Omega_{1}|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}$, we see that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda,$
which completes the proof. ∎
Remark 3.4.
Note that in Theorem 1.8, for $n\geq 4$,
$C_{2}(n-1)=\frac{n^{3/2}\pi}{C_{1}(n-1)},$
while for $n=2,3$
$\frac{n^{3/2}\pi}{C_{1}(n-1)}>C_{2}(n-1)=\begin{cases}2\sqrt{2}\pi,&n=2,\\ 9\sqrt{3},&n=3.\end{cases}$
This is because in the proof of two and three dimensional case, we use Corollary 2.4, which allows us to reduce the problem for large eigenvalues, in which cases by using Lemma 2.1 we have a better estimate than the uniform estimate from Theorem 2.3, see (3.1) and (3.5).
3.2 Three and higher dimensions
We next provide the proof for $n\geq 3$. Let $Q\subset{\mathbb{R}}^{n}$ be a unit cube. Similar to Lemma (3.2), we have
Lemma 3.5.
Let $Q\subset{\mathbb{R}}^{n}$ be a cube.
(i) For $n=3$, for any $\lambda>3\pi^{2}/\ell(Q)^{2}$, it holds that
$\mathcal{N}_{Q}^{D}(\lambda)\geq\frac{\ell(Q)^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{\omega(3)3^{3/2}\pi\ell(Q)^{2}}{(2\pi)^{3}}\lambda+1,$
and for all $\lambda>0$ that
$\mathcal{N}_{Q}^{D}(\lambda)\geq\frac{\ell(Q)^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{\omega(3)3^{3/2}\pi\ell(Q)^{2}}{(2\pi)^{3}}\lambda.$
(ii) For $n\geq 4$, for any $\lambda>n^{3}\pi^{2}/\ell(Q)^{2}$, it holds that
$\mathcal{N}_{Q}^{D}(\lambda)\geq\frac{\omega(n)\ell(Q)^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q)^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}+2^{n}\pi^{n/2},$
and for all $\lambda>0$ that
$\mathcal{N}_{Q}^{D}(\lambda)\geq\frac{\omega(n)\ell(Q)^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q)^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}.$
Proof.
(i) For $n=3$, we have
$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}+c^{2}),\,a,b,c\in\mathbb{N}\right\}.$
So the number $\mathcal{N}_{Q}^{D}(\lambda)$ is not less than the number of the unit cubes contained in the first quadrant intersecting the ball $B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{3})$, where $\sqrt{3}$ is the diameter of the unit cube. We therefore deduce that
$\mathcal{N}_{Q}^{D}(\lambda)$ $\geq\frac{1}{8}|B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{3})|=\frac{\pi}{6}\left(\ell(Q)\sqrt{\lambda}/\pi-\sqrt{3}\right)^{3}$
$\geq\frac{\ell(Q)^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{3\sqrt{3}\pi\ell(Q)^{2}}{6}\frac{\lambda}{\pi^{2}}+\frac{9\pi}{6}\frac{\ell(Q)\sqrt{\lambda}}{\pi}-\frac{3\sqrt{3}\pi}{6}$
$>\frac{\ell(Q)^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{3\sqrt{3}\ell(Q)^{2}}{6}\frac{\lambda}{\pi}+1,$
since $\sqrt{\lambda}>\pi\sqrt{3}/\ell(Q)$. For $\sqrt{\lambda}\leq\pi\sqrt{3}/\ell(Q)$, note that
$\frac{\ell(Q)^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{3\sqrt{3}\ell(Q)^{2}}{6}\frac{\lambda}{\pi}\leq 0\leq\mathcal{N}_{Q}^{D}(\lambda),$
which completes the proof.
(ii) For $n\geq 4$, we also have
$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}\sum_{i=1}^{n}a_{i}^{2},\,a_{i}\in\mathbb{N}\right\}.$
So the number $\mathcal{N}_{Q}^{D}(\lambda)$ is not less than the number of the unit cubes contained in the first quadrant intersecting the ball $B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{n})$, where $\sqrt{n}$ is the diameter of the unit cube. We therefore deduce that
$\mathcal{N}_{Q}^{D}(\lambda)$ $\geq\frac{1}{2^{n}}|B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{n})|=2^{-n}\omega(n)\left(\ell(Q)\sqrt{\lambda}/\pi-\sqrt{n}\right)^{n}$
$\geq 2^{-n}\omega(n)\sum_{i=0}^{n}C_{n}^{i}(\ell(Q)\lambda^{1/2}\pi^{-1})^{n-i}(-\sqrt{n})^{i}$
$>\frac{\omega(n)\ell(Q)^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q)^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}+2^{-n}\omega(n)\sum_{i=2}^{n}C_{n}^{i}(\ell(Q)\lambda^{1/2}\pi^{-1})^{n-i}(-\sqrt{n})^{i},$
where $C_{n}^{i}$ denotes the combinatoric number. So if
$\ell(Q)\lambda^{1/2}\pi^{-1}\geq n^{3/2},$
then we have
$\mathcal{N}_{Q}^{D}(\lambda)$ $\geq\frac{\omega(n)\ell(Q)^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q)^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}+\frac{\omega(n)n^{2}(n-1)}{3\times 2^{n}}(\ell(Q)\lambda^{1/2}\pi^{-1})^{n-2}n^{2}.$
Since
$\Gamma(1+\frac{n}{2})\leq\left(\frac{n+2}{2}\right)^{n/2},$
we deduce that
$\frac{\omega(n)}{2^{n}}\frac{n^{4}(n-1)}{3}(\ell(Q)\lambda^{1/2}\pi^{-1})^{n-2}\geq\frac{(\pi)^{n/2}}{2^{n}}\left(\frac{2}{n+2}\right)^{n/2}n^{\frac{3n}{2}+1}\geq 2^{n}\pi^{n/2},$
and hence,
$\mathcal{N}_{Q}^{D}(\lambda)$ $\geq\frac{\omega(n)\ell(Q)^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q)^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}+2^{n}\pi^{n/2}.$
For $\ell(Q)\lambda^{1/2}\pi^{-1}<n^{3/2},$ it holds that
$\frac{\omega(n)\ell(Q)^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q)^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}<0\leq\mathcal{N}_{Q}^{D}(\lambda),$
which completes the proof. ∎
Recall that by (1.9) and (2.4), it holds for $n\geq 4$ that
(3.3) $C_{1}(n-1)=\frac{3\pi\mathscr{B}(\frac{n-1}{2}-1,2)}{2\mathscr{B}(\frac{n-1}{2}-1,5/2)}C_{3}(\frac{n-1}{2})=\frac{3\pi\mathscr{B}(\frac{n-3}{2},2)}{2\mathscr{B}(\frac{n-3}{2},5/2)}C_{3}(\frac{n-1}{2}),$
where
$C_{3}(p)=\begin{cases}\frac{1}{9},&1<p\leq 2,\\ \frac{1}{8}\left(\frac{247}{256}\right)^{p-1},&p>2.\end{cases}$
Theorem 3.6.
Let $n\geq 3$ and suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture (e.g. ball, tiling domain, or our example in ${\mathbb{R}}^{n-1}$), $\Omega_{2}$ is an interval in ${\mathbb{R}}$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$ that is contained in $\Omega_{1}\times\Omega_{2}$. If $|\Omega_{1}|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $(\Omega_{1}\times\Omega_{2})\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where $C_{2}(n-1)$ is as in (1.10),
$C_{2}(n)=\begin{cases}9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
Proof.
Denote by $\Omega_{0}=\Omega_{1}\times\Omega_{2}$, $\Omega_{3}$ the union ${\cup_{k\in{\mathbb{N}}}Q_{k}}$, and $\Omega=\Omega_{0}\setminus\overline{\Omega_{3}}$. Let
$\mathscr{V}_{Q}:=\left|\cup_{k=1}^{\infty}Q_{k}\right|=\sum_{k=1}^{\infty}\ell(Q_{k})^{n}.$
Then
$\mathscr{V}_{Q}\leq\mathscr{S}_{Q}<\infty.$
Step 1. Let us first consider the case $n=3$. By Theorem 2.3 we have for $\lambda>0$ that
$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{{3\pi}}{16|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
For $\Omega$ we have by Corollary 2.4, for $k_{0}$ being the largest integer satisfying
$k_{0}\leq\frac{|\Omega|\omega(3)}{(2\pi)^{n}}\left[\left(\frac{\mathscr{V}_{Q}}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{3\pi}{16|\Omega_{2}|}\right]^{3}=\frac{9\pi|\Omega_{1}|^{3}|\Omega|}{2^{13}\mathscr{V}_{Q}^{3}},$
it holds for each $1\leq k\leq k_{0}$, the Pólya conjecture holds on $\Omega$ for $\lambda_{k}(\Omega)$, i.e.,
$k\leq\frac{|\Omega|\omega(3)}{(2\pi)^{3}}\lambda_{k}(\Omega)^{3/2}.$
Since $|\Omega_{1}|\geq 9\sqrt{3}\mathscr{S}_{Q}\geq 9\sqrt{3}\mathscr{V}_{Q}$ and $|\Omega_{2}|>1$, it holds that
$\lambda_{k_{0}}(\Omega)\geq\left(6\pi^{2}\frac{k_{0}}{|\Omega|}\right)^{2/3}>\left(6\pi^{2}\frac{\frac{9\pi|\Omega_{1}|^{3}|\Omega|}{2^{13}\mathscr{V}_{Q}^{3}}-1}{|\Omega|}\right)^{2/3}>\left(6\pi^{2}\frac{\frac{8\pi|\Omega_{1}|^{3}|\Omega|}{2^{13}\mathscr{V}_{Q}^{3}}}{|\Omega|}\right)^{2/3}>61>3\pi^{2}.$
Recall that Lemma 2.1 implies that for $\lambda>\lambda_{1}(\Omega_{2})$,
$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|\omega(2)}{(2\pi)^{2}}(\lambda-\lambda_{j}(\Omega_{2}))$
(3.4) $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right).$
Since $\lambda_{k}(\Omega)>61$, $|\Omega_{2}|>1$, we have that
$\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda_{k}(\Omega)}}{\pi}}\right)^{2}\geq\frac{4}{9},$
and hence,
(3.5) $\mathcal{N}^{D}_{\Omega_{0}}(\lambda)=\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{\pi}{3|\Omega_{2}|\sqrt{\lambda}}\right).$
For
$\lambda\geq\lambda_{k_{0}}(\Omega)>61,$
by the monotonicity of Dirichlet eigenvalues and the fact $|\Omega\cup\Omega_{3}|=|\Omega_{0}|$, we then have
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{{\pi}}{3|\Omega_{2}|\sqrt{\lambda}}\right).$
Note that by Lemma 3.5 (i),
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
$\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\left(\frac{\ell(Q_{j})^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{\sqrt{3}\ell(Q_{j})^{2}}{2}\frac{\lambda}{\pi}\right)$
$=\mathcal{N}^{D}_{\Omega}(\lambda)+\frac{\mathscr{V}_{Q}}{6\pi^{2}}\lambda^{3/2}-\frac{\sqrt{3}\mathscr{S}_{Q}}{2}\frac{\lambda}{\pi}.$
This together with (3.5) implies that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{\pi}{3|\Omega_{2}|\sqrt{\lambda}}\right)-\frac{\mathscr{V}_{Q}}{6\pi^{2}}\lambda^{3/2}+\frac{\sqrt{3}\mathscr{S}_{Q}}{2}\frac{\lambda}{\pi}$
$=\frac{|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}}{(2\pi)^{3}}\omega(3)\lambda^{3/2}-\lambda\left(\frac{|\Omega_{1}|}{18\pi}-\frac{\sqrt{3}\mathscr{S}_{Q}}{2\pi}\right)$
$=\frac{|\Omega|}{(2\pi)^{3}}\omega(3)\lambda^{3/2}-\lambda\left(\frac{|\Omega_{1}|}{18\pi}-\frac{\sqrt{3}\mathscr{S}_{Q}}{2\pi}\right).$
As $|\Omega_{1}|\geq 9\sqrt{3}\mathscr{S}_{Q}$, we see that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{(2\pi)^{3}}\omega(3)\lambda^{3/2},$
which completes the case $n=3$.
Step 2. Let us consider the case $n\geq 4$. The proof is the same as above, we only need to verify some constants.
For $\Omega_{0}$, by Theorem 2.3 we have for $\lambda>0$ that
$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
Moreover, for $\Omega$, by Corollary 2.4, we have for $k_{0}$ being the largest integer satisfying
$k_{0}$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n-1)}{|\Omega_{2}|}\right]^{n}=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(\frac{\mathscr{V}_{Q}}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n-1)}{|\Omega_{2}|}\right]^{n}$
(3.6) $=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\frac{(C_{1}(n-1)|\Omega_{1}|)^{n}}{\mathscr{V}_{Q}^{n}},$
the Pólya conjecture holds true for $1\leq k\leq k_{0}$ on $\Omega$, i.e.,
$\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}\geq k.$
So for $\lambda_{k_{0}}(\Omega)$, as $|\Omega_{1}|\geq n^{3/2}\pi\mathscr{S}_{Q}/C_{1}(n-1)\geq n^{3/2}\pi\mathscr{V}_{Q}/C_{1}(n-1)$, we have
$\lambda_{k_{0}}(\Omega)$ $\geq\left(\frac{(2\pi)^{n}}{|\Omega|\omega(n)}k_{0}\right)^{2/n}\geq\left(\frac{(2\pi)^{n}}{2|\Omega|\omega(n)}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\frac{(C_{1}(n-1)|\Omega_{1}|)^{n}}{\mathscr{V}_{Q}^{n}}\right)^{2/n}$
$\geq\frac{1}{2}\frac{(C_{1}(n-1)|\Omega_{1}|)^{2}}{\mathscr{V}_{Q}^{2}}>n^{3}\pi^{2}/2.$
For
$\lambda\geq\lambda_{k_{0}}(\Omega),$
by the monotonicity of Dirichlet eigenvalues and the fact $|\Omega\cup\Omega_{3}|=|\Omega_{0}|$, we then have
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|\Omega_{2}|\sqrt{\lambda}}\right).$
By Lemma 3.5 (ii),
$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
$\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\left(\frac{\omega(n)\ell(Q_{j})^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q_{j})^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\right)$
$=\mathcal{N}^{D}_{\Omega}(\lambda)+\frac{\omega(n)\mathscr{V}_{Q}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}.$
The above two inequalities imply that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|\Omega_{2}|\sqrt{\lambda}}\right)-\left(\frac{\omega(n)\mathscr{V}_{Q}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\right)$
$=\frac{|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}}{(2\pi)^{n}}\omega(n)\lambda^{n/2}-\frac{\lambda^{\frac{n-1}{2}}\omega(n)}{(2\pi)^{n}}\left(C_{1}(n-1)|\Omega_{1}|-n^{3/2}\pi\mathscr{S}_{Q}\right)$
$=\frac{|\Omega|}{(2\pi)^{n}}\omega(n)\lambda^{n/2}-\frac{\lambda^{\frac{n-1}{2}}\omega(n)}{(2\pi)^{n}}\left(C_{1}(n-1)|\Omega_{1}|-n^{3/2}\pi\mathscr{S}_{Q}\right).$
As $|\Omega_{1}|\geq n^{3/2}\pi\mathscr{S}_{Q}/C_{1}(n-1)=C_{2}(n-1)\mathscr{S}_{Q}$, we see that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{(2\pi)^{n}}\omega(n)\lambda^{n/2},$
which completes the proof. ∎
Proof of the case $n\geq 3$ in Theorem 1.8.
The result follows from Theorem 3.6 immediately. ∎
3.3 Some other class satisfying Pólya’s conjecture
We next provide some result regarding the case of general domains satisfying Pólya’s conjecture.
Theorem 3.7.
Let $\Omega_{1}\subset{\mathbb{R}}^{n}$, $n\geq 2$. Suppose that Pólya’s conjecture holds on $\Omega_{1}$, and $\{Q_{j}\}_{j\in{\mathbb{N}}}$ is of Admissible class contained in $\Omega_{1}$. Then it holds on $\Omega=\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}}$ that
(3.7) $k\leq\begin{cases}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},\,&\,k\leq 2,\\ \frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},\,&\,k>2.\end{cases}$
Proof.
The required inequalities followed from the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$), see [16].
Note that $\Omega\cup\cup_{j}Q_{j}\subsetneq\Omega_{1}$, and $|\Omega\cup\cup_{j}Q_{j}|=|\Omega_{1}|$. By Lemma 3.5 and Lemma 3.5 (ii), and the fact $\Omega$, $Q_{j}$ are mutually disjoint, we find that
$\mathcal{N}^{D}_{\Omega_{1}}(\lambda)$ $\geq\mathcal{N}^{D}_{\Omega\cup\cup_{j}Q_{j}}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
$\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\left(\frac{\omega(n)\ell(Q_{j})^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q_{j})^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\right)$
$=\mathcal{N}^{D}_{\Omega}(\lambda)+\frac{\omega(n)\mathscr{V}_{Q}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}},$
where
$\mathscr{V}_{Q}:=\left|\cup_{k=1}^{\infty}Q_{k}\right|=\sum_{k=1}^{\infty}\ell(Q_{k})^{n}.$
This together with the assumption that $\Omega_{1}$ satisfies Pólya’s conjecture implies that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{1}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}-\left(\frac{\omega(n)\mathscr{V}_{Q}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\right)$
$=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}.$
This is equivalent to that, for $k\geq 3$,
(3.8) $k\leq\frac{|\Omega|}{(2\pi)^{n}}\omega(n)\lambda_{k}^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}^{\frac{n-1}{2}},$
which completes the proof. ∎
Inspired by [15, Theorem 1.2], using Theorem 3.7 we have another class of domains satisfying Pólya’s conjecture.
Corollary 3.8.
Let $\Omega_{1}\subset{\mathbb{R}}^{n}$, $n\geq 2$. Suppose that Pólya’s conjecture holds on $\Omega_{1}$ (e.g., tiling domain, ball, or our example), and $\{Q_{j}\}_{j\in{\mathbb{N}}}$ is of Admissible class contained in $\Omega_{1}$. Then the Pólya conjecture holds on $\Omega=(\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}})\times\Omega_{2}$, whenever $\Omega_{2}\subset{\mathbb{R}}$ satisfies
$|\Omega_{2}|\leq\begin{cases}\frac{|\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}}|}{4\sqrt{2}\pi\mathscr{S}_{Q}},&\,n=2,\\ \frac{|\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}}|}{2\times 3^{5/2}\mathscr{S}_{Q}}\left(\frac{247}{256}\right)^{\frac{n-2}{2}},&\,n=3,\\ \frac{|\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}}|}{3\mathscr{S}_{Q}}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}\frac{\mathscr{B}(\frac{n-3}{2},2)}{2n^{3/2}\mathscr{B}(\frac{n-3}{2},5/2)},&\,n\geq 4.\end{cases}$
Proof.
The proof is similar to the proof of Theorem 2.3. Let $\Omega_{0}=\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}}$. Then for all $\lambda>\lambda_{1}(\Omega)$, where
(3.9) $\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{0})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{0}|^{2/n}\omega(n)^{2/n}}+\frac{\pi}{|\Omega_{2}|^{2}},$
we have via Theorem 3.7 that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $=\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\#\left\{k:\,\lambda_{k}(\Omega_{0})+\lambda_{j}(\Omega_{2})<\lambda\right\}=\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\#\left\{k:\,\lambda_{k}(\Omega_{0})<(\lambda-\lambda_{j}(\Omega_{2}))\right\}$
$\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\left[\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{2}))^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{2}))^{\frac{n-1}{2}}\right].$
For $n=2$, via Lemma 2.1 we find that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\left[\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{2}))^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{2}))^{\frac{n-1}{2}}\right]$
$\leq\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}\frac{2|\Omega_{2}|}{3\pi}\lambda^{3/2}\left(1-\frac{3\pi}{16|\Omega_{2}|\sqrt{\lambda}}\right)+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\frac{|\Omega_{2}|}{4}\lambda$
$=\frac{|\Omega_{0}||\Omega_{2}|}{6\pi^{2}}\lambda^{3/2}+\frac{\omega(n)\lambda}{(2\pi)^{n}}\left({n^{3/2}\pi\mathscr{S}_{Q}}\frac{|\Omega_{2}|}{4}-\frac{|\Omega_{0}|}{8}\right).$
For
$|\Omega_{2}|\leq\frac{|\Omega_{0}|}{4\sqrt{2}\pi\mathscr{S}_{Q}},$
we see that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{0}||\Omega_{2}|}{6\pi^{2}}\lambda^{3/2}.$
For $n=3$, we via Lemma 2.1 and Lemma 2.2 that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\left[\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{3}))^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{2}))^{\frac{n-1}{2}}\right]$
$\leq\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}\left(\frac{2|\Omega_{2}|}{3\pi\mathscr{B}(\frac{n-2}{2},2)}\lambda^{\frac{n+1}{2}}\mathscr{B}(\frac{n-2}{2},5/2)-\frac{1}{9}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}\lambda^{\frac{n}{2}}\right)$
$\quad+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\frac{2|\Omega_{2}|}{3\pi}\lambda^{3/2}$
$=\frac{|\Omega_{0}||\Omega_{2}|\omega(n+1)}{(2\pi)^{n+1}}\lambda^{\frac{n+1}{2}}+\frac{\omega(n)\lambda^{\frac{n}{2}}}{(2\pi)^{n}}\left(\frac{2n^{3/2}\mathscr{S}_{Q}|\Omega_{2}|}{3}-\frac{|\Omega_{0}|}{9}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}\right).$
The desired conclusion follows as soon as
$|\Omega_{2}|\leq\frac{|\Omega_{0}|}{2\times 3^{5/2}\mathscr{S}_{Q}}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}.$
For $n\geq 4$, we deduce via Lemma 2.2 that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\left[\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{2}))^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}(\lambda-\lambda_{j}(\Omega_{2}))^{\frac{n-1}{2}}\right]$
$\leq\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}\left(\frac{2|\Omega_{2}|}{3\pi\mathscr{B}(\frac{n-2}{2},2)}\lambda^{\frac{n+1}{2}}\mathscr{B}(\frac{n-2}{2},5/2)-\frac{1}{9}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}\lambda^{\frac{n}{2}}\right)$
$\quad+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\frac{2|\Omega_{3}|}{3\pi\mathscr{B}(\frac{n-3}{2},2)}\lambda^{\frac{n}{2}}\mathscr{B}(\frac{n-3}{2},5/2)$
$=\frac{|\Omega_{0}||\Omega_{2}|\omega(n+1)}{(2\pi)^{n+1}}\lambda^{\frac{n+1}{2}}+\frac{\omega(n)\lambda^{\frac{n}{2}}}{(2\pi)^{n}}\left(\frac{2n^{3/2}\mathscr{S}_{Q}|\Omega_{3}|\mathscr{B}(\frac{n-3}{2},5/2)}{3\mathscr{B}(\frac{n-3}{2},2)}-\frac{|\Omega_{0}|}{9}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}\right).$
Since
$|\Omega_{2}|\leq\frac{|\Omega_{0}|}{3}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}\frac{\mathscr{B}(\frac{n-3}{2},2)}{2n^{3/2}\mathscr{S}_{Q}\mathscr{B}(\frac{n-3}{2},5/2)},$
we see that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{0}||\Omega_{2}|\omega(n+1)}{(2\pi)^{n+1}}\lambda^{\frac{n+1}{2}},$
which completes the proof. ∎
4 Proof of the main results
In this section, we prove our main uniform estimates on Lipschitz domains. Let us recall the Whitney decomposition, see [14, Appendix J].
Proposition 4.1.
Let $\Omega$ be an open nonempty subset of ${{{\mathbb{R}}}^{n}}$. Then there exists a family of dyadic cubes $\{Q_{i}\}_{j}$ such that
(a) $\cup_{j}Q_{j}=\Omega$ and the $Q_{j}^{\prime}s$ have disjoint interiors.
(b) $\sqrt{n}\ell(Q_{j})\leq\mathrm{dist}(Q_{j},\Omega^{c})\leq 4\sqrt{n}\ell(Q_{j}).$
(c) If the boundaries of $Q_{j}$ and $Q_{k}$ touch, then
$\frac{1}{4}\leq\frac{\ell(Q_{j})}{\ell(Q_{k})}\leq 4.$
(d) For a given $Q_{j}$ there exist at most $12^{n}$ $Q_{k}$’s that touch it.
We first provide the proof for the estimate on the remainder of Weyl’s law.
Proof of Theorem 1.5.
Step 1. Upper bound for Dirichlet eigenvalues. Let $\{Q_{j}\}_{j\in{\mathbb{N}}}$ be a sequence of dyadic cubes that forms a Whitney decomposition of $\Omega$ by Proposition 4.1. Suppose that $B(x_{0},r_{in})$ is the largest inner ball contained in $\Omega$.
Let $Q_{l}$ be a cube from $\{Q_{j}\}_{j}$ such that $x_{0}\in\overline{Q_{l}}$. Then it holds that
$\sqrt{n}\ell(Q_{l})\leq\mathrm{dist}(Q_{l},\Omega^{c})\leq r_{in}\leq 5\sqrt{n}\ell(Q_{l}).$
If there is some other $Q_{i}$, with $\ell(Q_{i})>\ell(Q_{l})$, then
$\sqrt{n}\ell(Q_{i})\leq r_{in}\leq 5\sqrt{n}\ell(Q_{l}).$
Let $2^{-k_{0}}$ be the side length of the largest cube in $\{Q_{j}\}_{j}$. So it holds that
$\frac{r_{in}}{5\sqrt{n}}\leq 2^{-k_{0}}=\sup_{Q_{j}}\{\ell(Q_{j})\}\leq\frac{r_{in}}{\sqrt{n}}.$
For $0<\epsilon<r_{in}$, let
$\Omega_{\epsilon}=\{x\in\Omega:\,\mathrm{dist}(x,\partial\Omega)<\epsilon\}.$
Let $k_{\epsilon}\in\mathbb{Z}$ be such that
$5\sqrt{n}2^{-k_{\epsilon}}\leq\epsilon<5\sqrt{n}2^{-k_{\epsilon}+1}.$
Then $k_{\epsilon}<k_{0}$.
By Lemma 3.2 and Lemma 3.5, we have for any $\lambda>0$,
$\mathcal{N}_{\Omega}^{D}(\lambda)$ $\geq\sum_{\ell(Q_{j})>2^{-k_{\epsilon}}}\mathcal{N}_{Q_{j}}^{D}(\lambda)$
$=\sum_{m=k_{0}}^{k_{\epsilon}-1}\sum_{Q_{j}:\,\ell(Q_{j})=2^{-m}}\mathcal{N}_{Q_{j}}^{D}(\lambda)$
$\geq\sum_{m=k_{0}}^{k_{\epsilon}-1}\sum_{Q_{j}:\,\ell(Q_{j})=2^{-m}}\left(\frac{\omega(n)\ell(Q_{j})^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q_{j})^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\right).$
Note that all the cubes $\{Q_{j}\}_{j}$ with side length smaller than $2^{-k_{\epsilon}+1}$ is contained in $\Omega_{\epsilon}$. Moreover, the number of cubes of side length $2^{-j}$, $j\geq k_{0}$, is smaller than
$\frac{|\Omega_{5\sqrt{n}2^{-j}}|}{2^{-jn}}\leq\frac{C_{Lip}(\Omega)5\sqrt{n}2^{-j}|\partial\Omega|}{2^{-jn}}\leq 5\sqrt{n}2^{j(n-1)}C_{Lip}(\Omega)|\partial\Omega|,$
where we used the fact that
$|\Omega_{5\sqrt{n}2^{-j}}|=\left|\{x\in\Omega:\,\mathrm{dist}(x,\partial\Omega)<5\sqrt{n}2^{-j}\}\right|\leq 5\sqrt{n}2^{-j}C_{Lip}(\Omega)|\partial\Omega|,$
$C_{Lip}(\Omega)$ is defined as in (1.6).
We therefore deduce that
$\mathcal{N}_{\Omega}^{D}(\lambda)$ $\geq\frac{\omega(n)|\Omega\setminus\Omega_{\epsilon}|}{(2\pi)^{n}}\lambda^{n/2}-\sum_{m=k_{0}}^{k_{\epsilon}-1}5\sqrt{n}C_{Lip}(\Omega)|\partial\Omega|\frac{\omega(n)n^{3/2}\pi}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}$
$\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)|\Omega_{\epsilon}|}{(2\pi)^{n}}\lambda^{n/2}-5\sqrt{n}(k_{\epsilon}-k_{0})C_{Lip}(\Omega)|\partial\Omega|\frac{\omega(n)n^{3/2}\pi}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}.$
Note that
$|\Omega_{\epsilon}|=\left|\{x\in\Omega:\,\mathrm{dist}(x,\partial\Omega)<\epsilon\}\right|\leq C_{Lip}(\Omega)\epsilon|\partial\Omega|$
and
$2^{k_{\epsilon}-k_{0}}\leq{\frac{10\sqrt{n}}{\epsilon}}{\frac{r_{in}}{\sqrt{n}}}\leq\frac{10r_{in}}{\epsilon}.$
We therefore have
(4.1) $\mathcal{N}_{\Omega}^{D}(\lambda)$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\left(\epsilon\lambda^{1/2}+5n^{2}\pi\log_{2}\frac{10r_{in}}{\epsilon}\right).$
Let $\mathscr{R}=\mathscr{R}_{n-1}\times I_{min}$ be the minimal rectangle for $\Omega$. Since
$\lambda_{1}(\Omega)\geq\lambda_{1}(\mathscr{R})>\frac{\pi^{2}}{|I_{min}|^{2}},$
we see that
$\frac{\pi r_{in}}{|I_{min}|\sqrt{\lambda_{1}(\Omega)}}<r_{in}.$
So for each $k\in{\mathbb{N}}$ we can let
$\epsilon=\frac{\pi r_{in}}{|I_{min}|\sqrt{\lambda_{k}(\Omega)}}$
and deduce that
$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
The proof of Step 1 is complete.
Step 2. Lower bound for the Dirichlet eigenvalues. To prove the lower bound for the Dirichlet eigenvalues of $\Omega$, let us set $\Omega_{1}=\mathscr{R}\setminus\overline{\Omega}$. Obviously, $\Omega_{1}$ is also a Lipschitz domain. Moreover, it holds that
|$\left|\{x\in\Omega_{1}:\,\mathrm{dist}(x,\partial\Omega_{1})<\epsilon\}\right|$ $\leq\epsilon|\partial\mathscr{R}\cap\partial\Omega_{1}|+C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}\cap(\partial\Omega\setminus\partial\mathscr{R})|$
(4.2) $\leq C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}|.$
Recall also that $r_{in}(\Omega_{1})$ denotes the inner radius of $\Omega_{1}$.
Since $\Omega\cup\Omega_{1}\subset\mathscr{R}=\mathscr{R}_{n-1}\times I_{min}$ and $|\Omega\cup\Omega_{1}|=|\mathscr{R}|$, we have via the monotonicity of Dirichlet eigenvalues for any $\lambda>0$ that
$\mathcal{N}_{\Omega\cup\Omega_{1}}^{D}(\lambda)=\mathcal{N}_{\Omega}^{D}(\lambda)+\mathcal{N}_{\Omega_{1}}^{D}(\lambda)\leq\mathcal{N}_{\mathscr{R}}^{D}(\lambda).$
Applying Step 1 to $\Omega_{1}$, (4) and Theorem 2.3, we find that
$\mathcal{N}_{\Omega}^{D}(\lambda)$ $\leq\mathcal{N}_{\mathscr{R}}^{D}(\lambda)-\mathcal{N}_{\Omega_{1}}^{D}(\lambda)$
$\leq\frac{|\mathscr{R}_{n-1}||I_{min}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|I_{min}|\sqrt{\lambda}}\right)_{+}$
$\quad-\frac{\omega(n)|\Omega_{1}|}{(2\pi)^{n}}\lambda^{n/2}+\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega_{1}|}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\left(\epsilon\lambda^{1/2}+5n^{2}\pi\log_{2}\frac{10r_{in}(\Omega_{1})}{\epsilon}\right),$
for $0<\epsilon<r_{in}(\Omega_{1})$.
Since
$\lambda_{1}(\Omega)\geq\lambda_{1}(\mathscr{R})>\frac{\pi^{2}}{|I_{min}|^{2}},$
we see that
$\frac{\pi r_{in}(\Omega_{1})}{|I_{min}|\sqrt{\lambda_{1}(\Omega)}}<r_{in}(\Omega_{1}).$
So for each $k\in{\mathbb{N}}$ we can let
$\epsilon=\frac{\pi r_{in}(\Omega_{1})}{|I_{min}|\sqrt{\lambda_{k}(\Omega)}}$
and get that
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(C_{Lip}(\Omega)|\partial\Omega_{1}|\left(\frac{\pi r_{in}(\Omega_{1})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right).$
The proof is complete. ∎
Remark 4.2.
If $\Omega\subset{{{\mathbb{R}}}^{n}}$ is a convex domain, then obviously
$|\Omega_{\epsilon}|=\left|\{x\in\Omega:\,\mathrm{dist}(x,\partial\Omega)<\epsilon\}\right|\leq\epsilon|\partial\Omega|,$
i.e., $C_{Lip}(\Omega)=1$ in Step 1 above.
For Step 2, when dealing with the complementary set $\mathscr{R}\setminus\overline{\Omega}$, one can carry out a Whitney decomposition of the domain $\mathscr{R}\setminus\overline{\Omega}$, w.r.t. to the boundary $\partial\Omega$, see Figure 3.
Figure 3: Whitney decomposition for the complementary of a convex set
The convexity of $\mathscr{R}$ and $\Omega$ would imply that
|$\left|\{x\in\mathscr{R}\setminus\overline{\Omega}:\,\mathrm{dist}(x,\partial\Omega)<\epsilon\}\right|\leq|\partial\mathscr{R}\setminus\partial\Omega|\epsilon.$
Applying this estimate to Step 2 above gives that
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(|\partial\mathscr{R}\setminus\partial\Omega|\left(\frac{\pi r_{in}(\Omega_{1})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right).$
Obviously, |Ω||($|\partial\mathscr{R}\setminus\partial\Omega|\leq|\partial(\mathscr{R}\setminus\overline{\Omega})|$. The number $\Lambda(\epsilon,\Omega)$ from Theorem 1.3 can be taken smaller than (1.7) as
$\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}-\frac{C_{1}(n-1)}{2n}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon,$
where $C_{1}(n-1)$ is as in (1.9), see the following proof of Theorem 1.3.
We now prove the $\epsilon$-loss version of Pólya’s conjecture for large eigenvalues.
Proof of Theorem 1.3.
Let $\Omega_{1}=\mathscr{R}\setminus\overline{\Omega}$. By Theorem 1.5, it holds for each $k\in{\mathbb{N}}$ that
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\Omega)|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right)$
$\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}\left(1+\frac{C_{Lip}(\Omega)(2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|)}{\sqrt{\lambda_{k}}|\Omega|}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)\right).$
Taking derivative of the function
$f(\Lambda):=\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{\sqrt{\Lambda}|\Omega|}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda}}{\pi}\right),$
we see that
$f^{\prime}(\Lambda)$ $=-\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{2\Lambda^{3/2}|\Omega|}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda}}{\pi}\right)$
$\quad\quad+\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{\sqrt{\Lambda}|\Omega|}\frac{5n^{2}\pi\log_{2}e}{2\Lambda}.$
We have for each $\Lambda\geq\frac{1}{\mathrm{width}(\Omega)^{2}}$ that
$\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda}}{\pi}\geq\frac{10}{\pi}>e,$
and therefore $f(\Lambda)$ is monotone decreasing on $[\frac{1}{\mathrm{width}(\Omega)^{2}},\infty)$.
Note that for $\Lambda=\frac{1}{\mathrm{width}(\Omega)^{2}}$,
$f(\Lambda)=\frac{(2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|)\mathrm{width}(\Omega)}{|\Omega|}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10}{\pi}\right)>2n,$
$\lim_{\Lambda\to\infty}f(\Lambda)=0,$
and $C_{Lip}(\Omega)\geq 1$. So for any $\epsilon\in(0,1)$, we can let $\Lambda>\frac{1}{\mathrm{width}(\Omega)^{2}}$ be such that
$C_{Lip}(\Omega)\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{|\Omega|\sqrt{\Lambda}}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda}}{\pi}\right)=\epsilon,$
and deduce for any $\lambda_{k}(\Omega)\geq\Lambda$ that
$k$ $\leq(1+\epsilon)\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
The proof is complete. ∎
For open sets, we also have a uniform estimate. Let us remark that, since the boundary of an open set may be complicate, in general one can not have a uniform estimate in term of its boundary as in Theorem 1.5. Note that $C_{1}(n-1)$ is defined via (1.9).
Theorem 4.3.
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded open domain, $n\geq 2$, let $Q$ be a cube that contains $\Omega$. For any $20<\epsilon<\ell(Q)/$ 2, it holds for all $k\in{\mathbb{N}}$ that
(4.3) $k\leq\frac{|\Omega_{\epsilon}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+\frac{\omega(n)\lambda^{\frac{n-1}{2}}}{(2\pi)^{n}}\left(2n^{2}\pi\frac{|Q\setminus\Omega_{\epsilon}|}{\epsilon}-C_{1}(n-1)\ell(Q)^{n-1}\right).$
where $\Omega_{\epsilon}=\{x\in Q:\,\mathrm{dist}(x,\Omega)<\epsilon\}$. In particular, one can take $Q$ to be of side length $\mathrm{diam}(\Omega)$.
Proof.
Since $\epsilon<\ell(Q)/2$, we can divide $\mathscr{R}$ into the union of cubes $Q_{j}$, which are cubes of side length $2^{-k}\ell(Q)$, where $k\in\mathbb{Z}_{+}$ is such that
(4.4) $\epsilon/2<2^{-k}\ell(Q)\sqrt{n}\leq\epsilon.$
Denote by $\tilde{Q}$ the union of all $Q_{j}$, such that $Q_{j}$ entirely locates in $Q\setminus\overline{\Omega}$. Then the number of $Q_{j}$ that composes $\tilde{Q}$ is controlled by
#$\#\{Q_{j}:\,Q_{j}\subset\tilde{Q}\}\leq\frac{|Q\setminus\Omega|}{2^{-kn}\ell(Q)^{n}}.$
On the other hand, $\Omega\subset\mathscr{R}\setminus\tilde{Q}$, which is further contained in
$\{x\in Q:\,\mathrm{dist}(x,\Omega)<2^{-k}\ell(Q)\sqrt{n}\}\subset\Omega_{\epsilon}.$
By Theorem 2.3, we have
$\mathcal{N}^{D}_{Q}(\lambda)$ $\leq\frac{|Q|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{\ell(Q)\sqrt{\lambda}}\right)_{+}.$
This together with the monotonicity of Dirichlet eigenvalues implies that
$\frac{|Q|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{\ell(Q)\sqrt{\lambda}}\right)_{+}$ $\geq\mathcal{N}^{D}_{Q}(\lambda)\geq\mathcal{N}^{D}_{Q\setminus\tilde{Q}}(\lambda)+\mathcal{N}^{D}_{\tilde{Q}}(\lambda)$
(4.5) $\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\frac{|Q\setminus\Omega_{\epsilon}|}{2^{-kn}\ell(Q)^{n}}\mathcal{N}^{D}_{Q_{j}}(\lambda).$
Since for each $Q_{j}$, by Lemma 3.2 and Lemma 3.5, it holds that
$\mathcal{N}^{D}_{Q_{j}}(\lambda)\geq\frac{\omega(n)2^{-kn}\ell(Q)^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{n^{3/2}\pi\omega(n)2^{-k(n-1)}\ell(Q)^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}},$
we see that
$\frac{|Q\setminus\Omega_{\epsilon}|}{2^{-kn}|I_{min}|^{n}}\mathcal{N}^{D}_{Q_{j}}(\lambda)$ $\geq\frac{\omega(n)|Q\setminus\Omega_{\epsilon}|}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)\lambda^{\frac{n-1}{2}}}{(2\pi)^{n}}\frac{n^{3/2}\pi|Q\setminus\Omega_{\epsilon}|}{2^{-k}|I_{min}|}.$
This, together with (4.4), (4), implies that for
$\lambda>\lambda_{1}(\Omega)\geq\lambda_{1}(Q),$
it holds that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|Q|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{\ell(Q)\sqrt{\lambda}}\right)-\frac{\omega(n)|Q\setminus\Omega_{\epsilon}|}{(2\pi)^{n}}\lambda^{n/2}+\frac{\omega(n)\lambda^{\frac{n-1}{2}}}{(2\pi)^{n}}\frac{n^{3/2}\pi|Q\setminus\Omega_{\epsilon}|}{2^{-k}\ell(Q)}$
$=\frac{|\Omega_{\epsilon}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+\frac{\omega(n)\lambda^{\frac{n-1}{2}}}{(2\pi)^{n}}\left(2n^{2}\pi\frac{|Q\setminus\Omega_{\epsilon}|}{\epsilon}-C_{1}(n-1)\ell(Q)^{n-1}\right).$
The proof is complete. ∎
Acknowledgments
R. Jiang wishes to thank Prof. Xiao Zhong for explaining to him Pólya’s conjecture and helpful discussions at University of Jyväskylä around 2010. R. Jiang was partially supported by NNSF of China (12471094 & 11922114), F.H. Lin was in part supported by the National Science Foundation Grant DMS2247773 and DMS1955249.

### v2

@@
 where $\Lambda(\epsilon,\Omega)$ is given explicitly.
-This reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem.
-This estimate is based on a uniform estimate on the remainder of the Weyl law on Lipschitz domains, which appears to be the first quantitative estimate for the remainder of Weyl’s law since Weyl’s seminal work in the year 1911.
-We also provide in all dimensions $n\geq 2$, a class of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
+This estimate is based on a uniform estimate on the remainder of the Weyl law with explicit constants so that it reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem.
+Our arguments in deriving such uniform estimates yield also, in all dimensions $n\geq 2$, classes of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
 1 Introduction
@@
 The Weyl law is also important in the longstanding question “can one hear about the shape of a cavity”, see Lord Rayleigh [30] and Kac [23].
-Moreover, the Weyl law is even related to the number theory, Prime number theorem, Riemann hypothesis, via Connes’ trace formula, see Connes [4], Fathizadeh-Khalkhali [11] and the beamer by Khalkhali [22].
+Moreover, the Weyl law maybe related to the prime number theorem, via Connes’ trace formula, see Connes [4], Fathizadeh-Khalkhali [13] and the beamer by Khalkhali [24].
 Let us denote the remainder of Weyl’s law as
@@
 Subsequent to Weyl’s seminal work, Courant [5], Courant-Hilbert [6], Hörmander [19, 20], Seeley [35, 36] and many others (see Ivrii [22] for comprehensive description on related history) made contributions to improve the estimate on the remainder.
-Up to present, the best estimate is as $O(\lambda^{\frac{n-1}{2}})$ on piecewise smooth domains by Seeley [32, 33], see also Hörmander [18].
 Weyl also conjectured a sharper asymptotic behavior of Dirichlet eigenvalue states for domains with piecewise smooth boundary that
@@
 The conjecture has been proved by Ivrii [21] under the condition that the set of all periodic geodesic billiards in $\Omega$ has measure zero, which was proved by Safarov-Vassiliev [34] for convex analytic domains and polygons.
-However, none of the above mentioned results gave quantitative estimate for the remainder $R_{\Omega}(\lambda)$ (that is, these results states that certain asymptotic estimates valid for large $\lambda$’s without quantified the sizes of such $\lambda$’s), and consequently they do not give any quantitative estimate on eigenvalues.
-From both mathematical and physical point of view, it is rather natural to ask the following question.
+However, the above mentioned results for the remainder $R_{\Omega}(\lambda)$ state that certain asymptotic estimates valid only for large $\lambda$’s without quantifying the sizes of such $\lambda$’s, and consequently they do not give quantitative estimate on eigenvalues.
+[^1] After the first version of the present article was posted on arXiv, R.
+Frank and S.
+Larson has kindly informed us recent works [9] and [12], in which several uniform estimates for the remainder were already established, see the added remark at the end of the introduction.
+It is rather natural to ask the following question.
 Question 1.1.
 Is it possible to give a uniform and quantitative estimate for the remainder $R_{\Omega}(\lambda)$ of Weyl’s law?
-From the mathematical point of view, uniform and quantitative estimates are always more interesting. In fact, in 1954, Pólya [29] conjectured that, it should hold on arbitrary bounded open domain that
+In 1954, Pólya [32] conjectured that, it should hold on arbitrary bounded open domain that
 $k\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}^{n/2},\ \forall\,k\in{\mathbb{N}}.$
@@
 Filonov et al.
-[7] and [8] verified, via a method from number theory, Pólya’s conjecture on balls for all $n\geq 2$, and on annuli for $n=2$.
+[7] and [8] verified, via number theoretic arguments, Pólya’s conjecture on balls for all $n\geq 2$, and on annuli for $n=2$.
 For general domains $\Omega$, it was known that Pólya’s conjecture holds for the first two Dirichlet eigenvalues see [18] for the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$).
@@
 Note that the estimate (1.5) and its improvements are still a distance away from Pólya’s conjecture.
-Our main purpose of this paper is to prove Pólya’s conjecture up to an $\epsilon$-loss, and to give a solution to Question 1.1, on Lipschitz domains.
-1.2 Main results
-Let us first give a definition for Pólya’s conjecture up to an $\epsilon$-loss.
+Our main purpose of this paper is to reduce the $\epsilon$-loss version of Pólya’s conjecture to a computational problem, and to give a solution to Question 1.1 with explicit constants, on Lipschitz domains.
+Moreover, we shall also provide an approach to the Pólya conjecture that might be applied to general smooth domains, and provide in all dimensions, $n\geq 2$, a class of domains that satisfy Pólya’s conjecture.
+1.2 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
+Let us start from a definition for Pólya’s conjecture up to an $\epsilon$-loss.
 Definition 1.2.
@@
 $k$ $\leq(1+\epsilon)\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
-Noting that $2\leq\left(\frac{n+2}{n}\right)^{n/2}\to e$ as $2\leq n\to\infty$, the $\epsilon$-loss version of Pólya’s conjecture improves substantially (1.5) and its improvements on lower order terms, see [2, 12, 23, 28].
+Noting that $2\leq\left(\frac{n+2}{n}\right)^{n/2}\to e$ as $2\leq n\to\infty$, the $\epsilon$-loss version of Pólya’s conjecture improves substantially (1.5) (and its improvements on lower order terms, see [2, 14, 25, 31]).
 For a bounded open domain $\Omega$, we define its width $\mathrm{width}(\Omega)$ as the minimum distance between two parallel hyperplanes that contain $\Omega$ inside.
@@
 (1.6) |$\left|\{x\in{{{\mathbb{R}}}^{n}}:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)\leq\epsilon\}\right|\leq C_{Lip}(\Omega)\epsilon|\partial\Omega|.$
-Our first main result is as following.
+We have the following Pólya’s conjecture up to $\epsilon$-loss.
 Theorem 1.3.
@@
 (1.7) $\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}C_{Lip}(\Omega)\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
-Obviously, for a given $\epsilon\in(0,1)$, and a concrete Lipschitz domain $\Omega$, $\Lambda(\epsilon,\Omega)$ is uniquely determined by the equation (1.7).
-Hence, to show the $\epsilon$-loss version of Pólya’s conjecture, Theorem 1.3 reduces the problem to a computational problem for eigenvalues $\lambda_{k}(\Omega)<\Lambda(\epsilon,\Omega)$.
+Moreover, if $\Omega$ is convex, then $\Lambda(\epsilon,\Omega)$ can be taken smaller as
+(1.8) $\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
+Obviously, for a given $\epsilon\in(0,1)$, and a concrete Lipschitz (resp.
+convex) domain $\Omega$, $\Lambda(\epsilon,\Omega)$ is uniquely determined by the equation (1.7) (resp.
+(1.8) for convex domains).
+It can be seen from the above estimate that $\Omega$ needs only to be a bounded set of finite perimeter.
+Also, to show the $\epsilon$-loss version of Pólya’s conjecture, Theorem 1.3 reduces the problem to a computational problem for eigenvalues $\lambda_{k}(\Omega)<\Lambda(\epsilon,\Omega)$.
 Theorem 1.3 will follow from our following uniform estimate for the remainder of the Weyl’s law. Let us start from the following definition.
@@
 We say that $\mathscr{R}=I_{1}\times I_{2}\times\cdots\times I_{n}=\mathscr{R}_{n-1}\times I_{min}$ is the minimal admissible rectangle for $\Omega$, if $I_{min}=I_{n}=\mathrm{width}(\Omega)$, and $\mathscr{R}$ contains $\Omega$, and for each $i=1,\cdots,n$, for any $\tilde{I}_{i}$ with $|\tilde{I}_{i}|<|I_{i}|$, $I_{1}\times\cdots\tilde{I}_{i}\times\cdots\times I_{n}$ does not contain $\Omega$ inside.
-We have the following uniform estimate on Lipschitz domains, which gives a solution to Question 1.1.
-Moreover, our uniform estimate appears to be the first quantitative estimate for the remainder of Weyl’s law since Weyl’s seminal work [34] in the year 1911, see also Courant [5], Courant-Hilbert [6], Hörmander [18], Ivrii [19], Seeley [32, 33] etc for asymptotic results.
+We have the following uniform estimate on Lipschitz domains.
 In what follows, for a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$.
@@
 Moreover, it holds for each $k\in{\mathbb{N}}$ that
-$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
+(1.10) $k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
 Above,
 (1.11) $C_{1}(n)=\begin{cases}\frac{2}{3}&n=1,\\ \frac{3\pi}{16}&n=2,\\ \frac{\pi\mathscr{B}(\frac{n}{2}-1,2)}{6\mathscr{B}(\frac{n}{2}-1,5/2)}&n=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n}{2}-1,2)}{16\mathscr{B}(\frac{n}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
-Remark 1.6.
-If $\Omega\subset{{{\mathbb{R}}}^{n}}$ is a convex set, then one can take $C_{Lip}(\Omega)$ to be 1 in the above two theorems.
-In fact, one can have better estimates than (1.7) and (1.8).
-See Remark (4.2) below.
 Obviously, both quantities $\frac{r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}$ and $\frac{r_{in}(\Omega)}{|I_{min}|}$ are not larger than $1/2$, and therefore can be replaced by $1/2$.
 Moreover, note that if $|\partial(\mathscr{R}\setminus\overline{\Omega})|$ is small compared to $|\mathscr{R}_{n-1}|$, we then have that Pólya’s conjecture holds for the first some eigenvalues.
-Our quantitative estimate is consistent to the asymptotic estimates established by Courant [5] and Courant-Hilbert [6], which is, up to present, best possible in the case of domains having only Lipschitz continuous boundary.
+Our quantitative estimate is consistent to the asymptotic estimates established by Courant [5] and Courant-Hilbert [6].
+For convex sets, we have a clean and better estimate as following.
+Corollary 1.6.
+Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded convex domain, $n\geq 2$.
+Let $\mathscr{R}=\mathscr{R}_{n-1}\times I_{min}$ be the minimal Admissible rectangle for $\Omega$.
+It holds for all $k\in{\mathbb{N}}$ that
+$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
+$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(|\partial\mathscr{R}\setminus\partial\Omega|\left(\frac{\pi r_{in}(\mathscr{R}\setminus{\overline{\Omega}})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right),$
+and
+$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
 Let us remark that the logarithm term in (1.9) comes from the Whitney decomposition for $\mathscr{R}\setminus\overline{\Omega}$, see the proof of Theorem 1.5 in Section 4 below.
@@
 If $\mathscr{R}\setminus\overline{\Omega}$ admits a better decomposition, such as a union of up to infinite many cubes with finite surface, we then can prove full Pólya’s conjecture, see discussions in the next subsection.
-1.3 Pólya’s conjecture on a class of domains
+1.3 An approach to Pólya’s conjecture and concrete examples
 Using the strategy developed above, we can provide in all dimensions $n\geq 2$, a class of domains which may have rather irregular shape or boundary but satisfy Pólya’s conjecture.
+Using Seeley’s asymptotic estimates [35, 36] (see also Hörmander [20]), we first provide an approach to Pólya’s conjecture that might be applied to general smooth domains.
+By a rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$, we mean $\mathscr{R}_{n-1}\subset{\mathbb{R}}^{n-1}$ is a rectangle, and $I\subset{\mathbb{R}}$ is an open interval.
+Theorem 1.7.
+Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$.
+Assume that $\Omega$ is a piecewise smooth domain contained in $\mathscr{R}$.
+There exists $C_{P}=C(\Omega,n)$ such that if $|\mathscr{R}_{n-1}|\geq C_{P}$ then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\Omega}$.
 In what follows, by a cube $Q$ we always mean it is an open cube and denote by $\ell(Q)$ its side length.
@@
 We say that $\{Q_{k}\}_{k\in{\mathbb{N}}}$ is of Admissible class if $\mathscr{S}_{Q}<\infty$, and for any $k,j\in{\mathbb{N}}$, $Q_{k}\cap Q_{j}=\emptyset$ if $k\neq j$.
-By a rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$, we mean $\mathscr{R}_{n-1}\subset{\mathbb{R}}^{n-1}$ is a rectangle, and $I\subset{\mathbb{R}}$ is an open interval.
+We have a class of domains that satisfy Pólya’s conjecture in all dimensions.
 Theorem 1.9.
@@
 Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\mathscr{R}$.
-If $|\mathscr{R}_{n-1}|>C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
+If $|\mathscr{R}_{n-1}|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
 (1.12) $C_{2}(n)=\begin{cases}2\sqrt{2}\pi,&n=1,\\ 9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
-For $M\geq 2$, by $\tilde{\Omega}$ $M$-tiles $\Omega$ we mean there are $M$ isometries of $\tilde{\Omega}$, $\tilde{\Omega}_{1\leq k\leq M}$, with non-overlapped interior, such that $\cup_{1\leq k\leq M}\tilde{\Omega}_{k}\subset\Omega$ and $|\Omega|=M|\tilde{\Omega}|$.
+For $M\geq 2$, by $\tilde{\Omega}$ $M$-tiles $\Omega$ we mean there are $M$ isometries of $\tilde{\Omega}$, $\{\tilde{\Omega}_{k}\}_{1\leq k\leq M}$, with non-overlapped interior, such that $\cup_{1\leq k\leq M}\tilde{\Omega}_{k}\subset\Omega$ and $|\Omega|=M|\tilde{\Omega}|$.
 Corollary 1.10.
@@
 In fact, in higher dimensions, $n\geq 3$, the rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I$ can be replaced some more general product $\Omega\times I$, as soon as Pólya’s conjecture holds on $\Omega$, see Theorem 3.7 below.
-Although only a part of discussions in Sections 2 and 3 will be sufficient for the proofs of main results, the extended discussions are of independent interest and may be used for other purposes.
+Although only a part of discussions in Sections 2 and 3 will be sufficient for the proofs of Theorem 1.3 and Theorem 1.5, the extended discussions are of independent interest and they have other uses.
 For Theorem 1.5 we shall apply the Whitney decomposition of a Lipschitz domain $\Omega$, and using the lower bound for the counting function on cubes (see Lemma 3.3 and Lemma 3.6 below) along with an optimization to deduce a lower bound for the counting function $\mathcal{N}_{\Omega}^{D}(\lambda)$ on $\Omega$.
@@
 Theorem 1.3 then follows almost immediately from Theorem 1.5.
-Finally, let us point it out, our approach depends heavily on the min-max principle (monotonicity of Dirichlet eigenvalues), which is close to the original one by Courant [5] and Courant-Hilbert [6].
-However, the way we use the min-max principle is rather different from that in Courant [5] and Courant-Hilbert [6].
-By our choose of minimal Admissible rectangles, we do not even have to take care of the cubes which intersect the boundary of $\Omega$.
-The latter is the main issue for Courant [5] and Courant-Hilbert [6].
+Finally, let us point it out, our approach depends on a modified min-max argument, which is related to but different from the original ones by Courant [5] and Courant-Hilbert [6].
+By our choose of minimal Admissible rectangles, we do not have to take care of the term in Courant [5] and Courant-Hilbert [6] for cubes which intersect the boundary of $\Omega$.
+The latter is a difficult issue for Courant [5] and Courant-Hilbert [6].
 The paper is organized as follows.
 In Section 2, we study the refined estimates on product domains, and prove the main results Theorem 2.3 and Corollary 2.4.
-In Section 3, we provide the proof for the results on concrete examples of Pólya’s conjecture, Theorem 1.8.
+In Section 3, we provide the proofs for Theorem 1.7 and the results on concrete examples of Pólya’s conjecture, Theorem 1.9.
 Since in higher dimensions, $n\geq 3$, we can replace the rectangle $\mathscr{R}_{n-1}$ by general set satisfying Pólya’s conjecture, we shall split the proof into the case of two dimension, and three and higher dimensions where we prove a more general result, Theorem 3.7.
-We shall also provide a qualitative result on removing general piecewise smooth set, Theorem 3.1, and shall show some results regarding cubes of Admissible class from some more general class of domains, see Theorem 3.7 below.
+We shall also show some results regarding cubes of Admissible class from some more general class of domains, see Theorem 3.8 below.
 We shall employ Lemma 3.3 and Lemma 3.6 from Section 3 and Theorem 2.3 from Section 2 to prove the uniform estimates on Lipschitz sets, Theorem 1.5 and its corollary, Theorem 1.3.
@@
 For $C>0$ we use $\lfloor C\rfloor$ to denote the integer part of it.
+For $a\in{\mathbb{R}}$ we denote by $a_{+}$ the quantity $\max\{0,a\}$.
 For a cube $Q$ we denote by $\ell(Q)$ its side length.
@@
 Throughout the paper, most of constants are given explicitly, if not we shall point it out.
+Added in remark.
+After the first version of this article was posted on arXiv, we are kindly informed by Frank and Larson, that a similar version to our uniform remainder estimate for Weyl’s law (Theorem 1.5) already appears in [9, Theorem 3.14] and [12, Proposition 3.6], which both show that
+(1.13) $\left|\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|$ $\leq C(n)\Theta_{\Omega}\lambda^{\frac{n-1}{2}}\left(1+\log_{+}(r_{in}(\Omega)\sqrt{\lambda})\right),$
+where $C(n)$ is an implicit constant depending only on the dimension $n$,
+$\Theta_{\Omega}:=\sup_{\delta>0}\delta^{-1}|\left\{x\in\Omega:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)<\delta\right\}|<\infty,$
+$\Theta_{\Omega}$ plays the same role of $C_{Lip}(\Omega)|\partial\Omega|$ in our estimate.
+In particular, it appears from the proof of [9, Theorem 3.14] that the above constant $C(n)$ can also be calculated.
+The proof of [12, Proposition 3.6] uses Tauberian theory, while the proof of [9, Theorem 3.14] uses Courant’s method of Dirichlet-Neumann bracketing and some refined estimates for cubes intersecting the boundary.
+Our proof for estimates of the reminder of Weyl’s law (Theorem 1.5) uses the minimal Admissible rectangle (see Theorem 2.3 below for refined eigenvalue estimates on product domains), a subtracting argument that allows us to avoid the cubes intersecting the boundary, and then lower bound for the counting function of Dirichlet eigenvalues.
+Moreover, our method has an advantage that gives many “irregular shaped domains” that verify the Pólya conjecture (see Theorem 1.9, Remark 1.11 and Theorem 3.7 below).
 2 Refined eigenvalue estimates on product domains
@@
 Moreover, if $k_{\lambda}\neq\frac{|\Omega|}{\pi}\sqrt{\lambda}$, there is another triangle insider the 1/4 ball with vertices at $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$, $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$ and $(\frac{|\Omega|}{\pi}\sqrt{\lambda},0)$.
-See Figure 2111Picture drawn by Xinyi Jiang..
+See Figure $2$Picture drawn by Xinyi Jiang..
 Figure 2: The 1/2 sum is controlled by the 1/4 ball
@@
 i.e.,
-$\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}<\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}=\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
+$\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}<\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\leq\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
 Let us prove (2.3). Since $\Omega$ is connected, we have
@@
 In view of Laptev’s beautiful theory on product domains [26] (which has been further employed in Laptev-Weidl [27] for Schrödinger operator), the Aizenman-Lieb principle [1], we have the following refined eigenvalue estimates on product domains if one domain satisfies Pólya’s conjecture.
+We note that Larson [28] earlier had obtained similar improvements for higher dimensions ($n_{1}\geq 3$ below) for more general products but with less explicit constants.
 Note that any bounded domain $\Omega\subset{\mathbb{R}}$ satisfies Pólya’s conjecture, i.e., $\lambda_{k}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}}k^{2}$ for each $k\in{\mathbb{N}}$.
@@
 $k_{0}=\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n_{1})}{|\Omega_{2}|}\right]^{n}\right\rfloor.$
-Then for each $1\leq k\leq k_{0}$, the Pólya conjecture holds on $\Omega$ for $\lambda_{k}(\Omega)$, i.e.,
+Then for each $0\leq k\leq k_{0}$, the Pólya conjecture holds on $\Omega$ for $\lambda_{k}(\Omega)$, i.e.,
 $k\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
@@
 $k_{0}=\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n_{1})}{|\Omega_{2}|}\right]^{n}\right\rfloor\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{0}^{n/2}.$
-Let $1\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
+Let $0\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
 $k\leq\mathcal{N}^{D}_{\Omega}(\lambda_{k}(\Omega)+\epsilon)$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}(\lambda_{k}(\Omega)+\epsilon)^{n/2}.$
@@
 The proof is complete. ∎
+Obviously the above corollary only makes sense when $k_{0}\geq 3$.
 Let us consider $\Omega_{1}\subset{\mathbb{R}}^{2}$, with $|\Omega_{1}|=100$, $\Omega_{2}\subset{\mathbb{R}}$ with $|\Omega_{2}|=1$.
@@
 3 Pólya’s conjecture on a class of domains
-Let us start with proving the qualitative result regarding removing general piecewise smooth set.
+Let us start with proving the qualitative result regarding removing general piecewise smooth set. Theorem 1.7 follows immediately from the following result.
 Theorem 3.1.
@@
 Up to a scaling, we may assume that $|\Omega_{3}|=1$.
-It follows from Seeley [32, 33] (see also [18, Corollary 17.5.11 & (17.5.16)’]) that for some $\Lambda>0$, it holds for $\lambda>\Lambda$ that
+It follows from Seeley [35, 36] (see also [20, Corollary 17.5.11 & (17.5.16)’]) that for some $\Lambda>0$, it holds for $\lambda>\Lambda$ that
 (3.1) $\left|\mathcal{N}^{D}_{\Omega_{3}}(\lambda)-\frac{|\Omega_{3}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|\leq C(\Omega_{3})\lambda^{\frac{n-1}{2}}.$
@@
 The proof is complete. ∎
+Remark 3.2.
+Note that (3.1) can be stated for all $\lambda>0$ without referring to $\Lambda$, but generally with larger constant $C(\Omega_{3})$.
+We state in this way to show that the combined argument, that is for small eigenvalues using Corollary 2.4 while for large constant using the subtracting argument, will give better estimate, see Remark 3.5 below.
 We next prove the concrete examples of domains obtained by removing cubes, Theorem 1.9. Let us start with the two dimensional case.
@@
 Note that for a given $\lambda_{k}$, $\mathcal{N}_{Q}^{D}(\lambda_{k})$ is dominated by the number of unit cubes contained in the first quadrant intersecting the ball $B(0,\sqrt{\lambda_{k}}/\pi)$.
-So it is readily to see that for all $k\in{\mathbb{N}}$ that $\lambda_{k}\geq 4\pi k.$ In fact Theorem 2.3 gives a better result. We however shall need the opposite estimates.
+So it is readily to see that for all $k\in{\mathbb{N}}$ that $\lambda_{k}\geq 4\pi k.$ In fact Theorem 2.3 gives a better result.
+We however shall need the opposite estimates.
 Lemma 3.3.
@@
 $\geq\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q)}{2}\sqrt{\lambda}+\frac{\pi}{2}.$
-Note that for $\lambda\leq 2\pi^{2}/\ell(Q)^{2}$,
-$\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q)}{2}\sqrt{\lambda}<0<\mathcal{N}_{Q}^{D}(\lambda),$
+Noting that for $\lambda\leq 2\pi^{2}/\ell(Q)^{2}$,
+$\frac{\ell(Q)^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q)}{2}\sqrt{\lambda}\leq 0\leq\mathcal{N}_{Q}^{D}(\lambda),$
 the conclusion is obvious. The proof is complete. ∎
@@
 $\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{0}}(\lambda).$
-Moreover, by Lemma 3.2 [(ii)], we see that for $\lambda\geq\lambda_{k_{0}}(\Omega)$,
-$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}_{Q_{k}}(\lambda)$
+Moreover, by Lemma 3.3 (ii), we see that for $\lambda\geq\lambda_{k_{0}}(\Omega)$,
+$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{k}}(\lambda)$
 $\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right).$
@@
 3.2 Three and higher dimensions
-We next provide the proof for $n\geq 3$. Let $Q\subset{\mathbb{R}}^{n}$ be a unit cube. Similar to Lemma (3.2), we have
+We next provide the proof for $n\geq 3$. Similar to Lemma (3.3), we have
 Lemma 3.6.
@@
 $>\frac{\ell(Q)^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{3\sqrt{3}\ell(Q)^{2}}{6}\frac{\lambda}{\pi}+1,$
-since $\sqrt{\lambda}>\pi\sqrt{3}/\ell(Q)$. For $\sqrt{\lambda}\leq\pi\sqrt{3}/\ell(Q)$, note that
+for $\sqrt{\lambda}>\pi\sqrt{3}/\ell(Q)$. For $\sqrt{\lambda}\leq\pi\sqrt{3}/\ell(Q)$, note that
 $\frac{\ell(Q)^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{3\sqrt{3}\ell(Q)^{2}}{6}\frac{\lambda}{\pi}\leq 0\leq\mathcal{N}_{Q}^{D}(\lambda),$
@@
 $k\leq\frac{|\Omega|\omega(3)}{(2\pi)^{3}}\lambda_{k}(\Omega)^{3/2}.$
-Since $|\Omega_{1}|\geq 9\sqrt{3}\mathscr{S}_{Q}\geq 9\sqrt{3}\mathscr{V}_{Q}$ and $|\Omega_{2}|>1$, it holds that
+Since $|\Omega_{1}|\geq 9\sqrt{3}\mathscr{S}_{Q}\geq 9\sqrt{3}\mathscr{V}_{Q}$ and $|\Omega_{2}|\geq 1$, it holds that
 $\lambda_{k_{0}}(\Omega)\geq\left(6\pi^{2}\frac{k_{0}}{|\Omega|}\right)^{2/3}>\left(6\pi^{2}\frac{\frac{9\pi|\Omega_{1}|^{3}|\Omega|}{2^{13}\mathscr{V}_{Q}^{3}}-1}{|\Omega|}\right)^{2/3}>\left(6\pi^{2}\frac{\frac{8\pi|\Omega_{1}|^{3}|\Omega|}{2^{13}\mathscr{V}_{Q}^{3}}}{|\Omega|}\right)^{2/3}>61>3\pi^{2}.$
@@
 (3.4) $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right).$
-Since $\lambda_{k}(\Omega)>61$, $|\Omega_{2}|>1$, we have that
+Since $\lambda_{k}(\Omega)>61$, $|\Omega_{2}|\geq 1$, we have that
 $\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda_{k}(\Omega)}}{\pi}}\right)^{2}\geq\frac{4}{9},$
@@
 Note that by Lemma 3.6 (i),
-$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
+$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{j}}(\lambda)$
 $\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\left(\frac{\ell(Q_{j})^{3}}{6\pi^{2}}\lambda^{3/2}-\frac{\sqrt{3}\ell(Q_{j})^{2}}{2}\frac{\lambda}{\pi}\right)$
@@
 By Lemma 3.6 (ii),
-$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
+$\mathcal{N}^{D}_{\Omega\cup\Omega_{3}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{j}}(\lambda)$
 $\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\left(\frac{\omega(n)\ell(Q_{j})^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q_{j})^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\right)$
@@
 Proof.
-The required inequalities followed from the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$), see [16].
+The required inequalities for $k\leq 2$ followed from the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$), see [18].
+Let us consider $k\geq 3$.
 Note that $\Omega\cup\cup_{j}Q_{j}\subsetneq\Omega_{1}$, and $|\Omega\cup\cup_{j}Q_{j}|=|\Omega_{1}|$.
 By Lemma 3.6 and Lemma 3.6 (ii), and the fact $\Omega$, $Q_{j}$ are mutually disjoint, we find that
-$\mathcal{N}^{D}_{\Omega_{1}}(\lambda)$ $\geq\mathcal{N}^{D}_{\Omega\cup\cup_{j}Q_{j}}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}_{Q_{j}}(\lambda)$
+$\mathcal{N}^{D}_{\Omega_{1}}(\lambda)$ $\geq\mathcal{N}^{D}_{\Omega\cup\cup_{j}Q_{j}}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{j}}(\lambda)$
 $\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{j\in{\mathbb{N}}}\left(\frac{\omega(n)\ell(Q_{j})^{n}}{(2\pi)^{n}}\lambda^{n/2}-\frac{\omega(n)n^{3/2}\pi\ell(Q_{j})^{n-1}}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\right)$
@@
 which completes the proof. ∎
-Inspired by [15, Theorem 1.2], using Theorem 3.7 we have another class of domains satisfying Pólya’s conjecture.
+Inspired by [17, Theorem 1.2], using Theorem 3.8 we have another class of domains satisfying Pólya’s conjecture.
 Corollary 3.9.
@@
 which completes the proof. ∎
-4 Proof of the main results
-In this section, we prove our main uniform estimates on Lipschitz domains. Let us recall the Whitney decomposition, see [14, Appendix J].
+4 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
+In this section, we prove our main uniform estimates on Lipschitz domains. Let us recall the Whitney decomposition, see [16, Appendix J].
 Proposition 4.1.
@@
 $5\sqrt{n}2^{-k_{\epsilon}}\leq\epsilon<5\sqrt{n}2^{-k_{\epsilon}+1}.$
-Then $k_{\epsilon}<k_{0}$.
+Then $k_{\epsilon}>k_{0}$.
 By Lemma 3.3 and Lemma 3.6, we have for any $\lambda>0$,
@@
 Obviously, $\Omega_{1}$ is also a Lipschitz domain.
-Moreover, it holds that
+Moreover, by (1.6), it holds that
 |$\left|\{x\in\Omega_{1}:\,\mathrm{dist}(x,\partial\Omega_{1})<\epsilon\}\right|$ $\leq\epsilon|\partial\mathscr{R}\cap\partial\Omega_{1}|+C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}\cap(\partial\Omega\setminus\partial\mathscr{R})|$
@@
 The proof is complete. ∎
-Remark 4.2.
+We next sketch the proof of Corollary 1.6.
+Proof of Corollary 1.6.
 If $\Omega\subset{{{\mathbb{R}}}^{n}}$ is a convex domain, then obviously
@@
 $k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
-$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(|\partial\mathscr{R}\setminus\partial\Omega|\left(\frac{\pi r_{in}(\Omega_{1})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right).$
-Obviously, |Ω||($|\partial\mathscr{R}\setminus\partial\Omega|\leq|\partial(\mathscr{R}\setminus\overline{\Omega})|$.
-The number $\Lambda(\epsilon,\Omega)$ from Theorem 1.3 can be taken smaller than (1.7) as
-$\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}-\frac{C_{1}(n-1)}{2n}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon,$
-where $C_{1}(n-1)$ is as in (1.9), see the following proof of Theorem 1.3.
+$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(|\partial\mathscr{R}\setminus\partial\Omega|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right).$
+The proof is complete. ∎
 We now prove the $\epsilon$-loss version of Pólya’s conjecture for large eigenvalues.
@@
 $k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
-$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\Omega)|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right)$
+$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\overline{\Omega})|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right)$
 $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}\left(1+\frac{C_{Lip}(\Omega)(2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|)}{\sqrt{\lambda_{k}}|\Omega|}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)\right).$
@@
 $k$ $\leq(1+\epsilon)\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
-The proof is complete. ∎
+If $\Omega$ is convex, then applying Corollary 1.6 gives that
+$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}\left(1+\frac{2n\mathrm{diam}(\Omega)^{n-1}}{\sqrt{\lambda_{k}}|\Omega|}\left(\frac{\pi}{2}-\frac{C_{1}(n-1)}{2n}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\lambda_{k}}}{\pi}\right)\right).$
+The remaining proof is the same. The proof is complete. ∎
 For open sets, we also have a uniform estimate.
-Let us remark that, since the boundary of an open set may be complicate, in general one can not have a uniform estimate in term of its boundary as in Theorem 1.5.
+Let us remark that, since the boundary of an open set may be complicate, in general one can not have a uniform estimate in term of its boundary as in Theorem 1.5; see [9, Theorem 3.14] for similar but earlier result.
 Note that $C_{1}(n-1)$ is defined via (1.11).
@@
 Acknowledgments
+We wish to thank Prof.
+Rupert Frank and Prof.
+Simon Larson for bringing references [9, 12, 28] to us and several useful comments.
 R.

### v3

@@
 where $\Lambda(\epsilon,\Omega)$ is given explicitly.
-This estimate is based on a uniform estimate on the remainder of the Weyl law with explicit constants so that it reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem.
+This reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem.
+This estimate is based on quantitative estimates on the remainder of the Weyl law with explicit constants, which we give a new proof without using Neumann eigenvalues.
 Our arguments in deriving such uniform estimates yield also, in all dimensions $n\geq 2$, classes of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
+Another key observation is that on strip-tiling domains (and therefore any triangles for instance) one actually has better eigenvalue estimates than Pólya conjectured.
 1 Introduction
@@
 In 1911, Weyl [39] proved the famous Weyl’s law, which states that for any bounded open domain $\Omega\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$, and $\lambda\to\infty$,
-(1.1) $\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+o(\lambda^{n/2}),$
+$\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+o(\lambda^{n/2}),$ (1.1)
 where $\omega(n)$ denotes the volume of unit ball in ${{\mathbb{R}}}^{n}$}, $\mathcal{N}^{D}_{\Omega}(\lambda)$ is the counting function
@@
 Let us denote the remainder of Weyl’s law as
-(1.2) $R_{\Omega}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$
+$R_{\Omega}(\lambda)=\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$ (1.2)
 Subsequent to Weyl’s seminal work, Courant [5], Courant-Hilbert [6], Hörmander [19, 20], Seeley [37, 38] and many others (see Ivrii [22] for comprehensive description on related history) made contributions to improve the estimate on the remainder.
 Weyl also conjectured a sharper asymptotic behavior of Dirichlet eigenvalue states for domains with piecewise smooth boundary that
-(1.3) $\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}-\frac{|\partial\Omega|}{2^{n+1}\pi^{\frac{d-1}{2}}\Gamma(\frac{n+1}{2})}\lambda^{\frac{n-1}{2}}+o(\lambda^{\frac{n-1}{2}}),$
+$\mathcal{N}^{D}_{\Omega}(\lambda)=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}-\frac{|\partial\Omega|}{2^{n+1}\pi^{\frac{d-1}{2}}\Gamma(\frac{n+1}{2})}\lambda^{\frac{n-1}{2}}+o(\lambda^{\frac{n-1}{2}}),$ (1.3)
 here and in what follows, for any set $\Omega$, $|\partial\Omega|$ denotes the surface area of it, $\mathscr{B}(\cdot,\cdot)$ the Beta function and $\Gamma(\cdot)$ the Gamma function, as usual.
@@
 However, the above mentioned results for the remainder $R_{\Omega}(\lambda)$ state that certain asymptotic estimates valid only for large $\lambda$’s without quantifying the sizes of such $\lambda$’s, and consequently they do not give quantitative estimate on eigenvalues.
-[^1] After the first version of the present article was posted on arXiv, R.
-Frank and S.
-Larson has kindly informed us recent works [9] and [12], in which several uniform estimates for the remainder were already established, see the added remark at the end of the introduction.
 It is rather natural to ask the following question.
@@
 Is it possible to give a uniform and quantitative estimate for the remainder $R_{\Omega}(\lambda)$ of Weyl’s law?
+It was proved earlier by Netrusov and Safarov [32], that a quantitative bound up to an extra logarithm term of the remainder of Weyl’s law holds on Lipshcitz type domains, see also [9, 12].
+[^1] After the first version of the present article was posted on arXiv, R.
+Frank and S.
+Larson has kindly informed us recent works [9] and [12], X.
+He informed us the work [32], in which several uniform estimates for the remainder were already established.
 In 1954, Pólya [34] conjectured that, it should hold on arbitrary bounded open domain that
@@
 For general domains, Berezin [2] and Li-Yau [29] independently proved for arbitrary bounded open domains that
-(1.4) $\frac{|\Omega|^{2/n}\omega(n)^{2/n}}{(2\pi)^{2}}\sum_{i=1}^{k}\lambda_{i}\geq\frac{n}{n+2}k^{\frac{n+2}{n}},$
+$\frac{|\Omega|^{2/n}\omega(n)^{2/n}}{(2\pi)^{2}}\sum_{i=1}^{k}\lambda_{i}\geq\frac{n}{n+2}k^{\frac{n+2}{n}},$ (1.4)
 which implies that
-(1.5) $\left(\frac{n+2}{n}\right)^{n/2}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}^{n/2}\geq k,$
+$\left(\frac{n+2}{n}\right)^{n/2}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}^{n/2}\geq k,$ (1.5)
 see also Laptev [26], and [10, 11, 14, 25, 31] for further improvements on the lower order term and generalizations.
 Note that the estimate (1.5) and its improvements are still a distance away from Pólya’s conjecture.
-Our main purpose of this paper is to reduce the $\epsilon$-loss version of Pólya’s conjecture to a computational problem, and to give a solution to Question 1.1 with explicit constants, on Lipschitz domains.
-Moreover, we shall also provide an approach to the Pólya conjecture that might be applied to general smooth domains, and provide in all dimensions, $n\geq 2$, a class of domains that satisfy Pólya’s conjecture.
+Our main purpose of this paper is to establish an approach to Pólya’s conjecture on more general domains (than tiling domains and products).
+More precisely, given a Lipschitz domain $\Omega$, we can choose a larger domain $T$, so that by using the monotone property of Dirichlet eigenvalues, one has
+$\mathcal{N}^{D}_{\Omega}(\lambda)\leq\mathcal{N}^{D}_{T}(\lambda)-\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda).$
+The Pólya conjecture will follow if one can prove, on one side $T$ satisfies better eigenvalue estimates than Pólya conjectured, and one the other side, $\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda)$ has a reasonable good lower bound.
+On one side, we shall prove that strip-tiling domains and certain product domains satisfy better eigenvalue estimates than Pólya conjectured.
+On the other side, we shall prove on certain domains composed by cubes of different scales the eigenvalue counting functions do have a good lower bound.
+These enable us to provide in all dimensions, $n\geq 2$, a class of domains that satisfy Pólya’s conjecture.
+For general Lipschitz domains, we did not have a good control of lower bound of $\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda)$, and our method only yields a new proof of quantitative Courant type estimate for the remainder.
+Let us emphasize that quantitative Courant type estimate for the remainder has already obtained in [32] (see also [7]) via Courant’s method of Dirichlet-Neumann bracketing, our estimate (see Theorem 1.5 and Corollary 1.6 below) give slightly better constants.
+If one can further obtain better estimate on the lower bound of $\mathcal{N}^{D}_{T\setminus\overline{\Omega}}(\lambda)$, then one can combine the refined estimate on $\mathcal{N}^{D}_{T}(\lambda)$ (see Remark 2.5 below, which behaves better as $\lambda\to\infty$), to reduce Pólya’s conjecture to a computational problem.
+In this paper, we can only reduce the $\epsilon$-loss version of Pólya’s conjecture to a computational problem.
 1.2 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
@@
 For a bounded Lipschitz domain $\Omega$, we let $C_{Lip}(\Omega)\geq 1$ be the constant, uniquely determined by $\Omega$, such that for $\epsilon\in(0,\mathrm{width}(\Omega))$, such that
-(1.6) |$\left|\{x\in{{{\mathbb{R}}}^{n}}:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)\leq\epsilon\}\right|\leq C_{Lip}(\Omega)\epsilon|\partial\Omega|.$
+$\left|\{x\in{{{\mathbb{R}}}^{n}}:\,{\mathop{\mathrm{\,dist\,}}}(x,\partial\Omega)\leq\epsilon\}\right|\leq C_{Lip}(\Omega)\epsilon|\partial\Omega|.$ (1.6)
 We have the following Pólya’s conjecture up to $\epsilon$-loss.
 Theorem 1.3.
-Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded Lipschitz domain, $n\geq 2$. For any $10<\epsilon<$ 1, it holds for all $\lambda_{k}(\Omega)\geq\Lambda(\epsilon,\Omega)$ that
+Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded Lipschitz domain, $n\geq 2$. For any $0<\epsilon<1$, it holds for all $\lambda_{k}(\Omega)\geq\Lambda(\epsilon,\Omega)$ that
 $k$ $\leq(1+\epsilon)\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},$
 where $\Lambda(\epsilon,\Omega)\geq\mathrm{width}(\Omega)^{-2}$ is such that
-(1.7) $\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}C_{Lip}(\Omega)\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
+$\frac{2n\mathrm{diam}(\Omega)^{n-1}+|\partial\Omega|}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}C_{Lip}(\Omega)\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$ (1.7)
 Moreover, if $\Omega$ is convex, then $\Lambda(\epsilon,\Omega)$ can be taken smaller as
-(1.8) $\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
+$\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$ (1.8)
 Obviously, for a given $\epsilon\in(0,1)$, and a concrete Lipschitz (resp.
@@
 Also, to show the $\epsilon$-loss version of Pólya’s conjecture, Theorem 1.3 reduces the problem to a computational problem for eigenvalues $\lambda_{k}(\Omega)<\Lambda(\epsilon,\Omega)$.
-Theorem 1.3 will follow from our following uniform estimate for the remainder of the Weyl’s law. Let us start from the following definition.
+Theorem 1.3 follows directly from our following uniform estimate for the remainder of the Weyl’s law. Let us start from the following definition.
 Definition 1.4 (Minimal Admissible rectangle).
@@
 We have the following uniform estimate on Lipschitz domains.
+Let us emphasize once more the following Courant type quantitative estimate has been obtained already in [32, Theorem 1.8] via Courant’s method of Dirichlet-Neumann bracketing.
+Our method for the proof of Theorem 1.5 and Corollary 1.6 below does not use Neumann eigenvalue and give better constant estimates.
 In what follows, for a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$.
@@
 where
-(1.9) $C(n,\Omega,\lambda_{k}(\Omega))=C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\overline{\Omega})|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|.$
+$C(n,\Omega,\lambda_{k}(\Omega))=C_{Lip}(\Omega)|\partial(\mathscr{R}\setminus\overline{\Omega})|\left(\frac{\pi r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|.$ (1.9)
 Moreover, it holds for each $k\in{\mathbb{N}}$ that
-(1.10) $k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
+$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$ (1.10)
 Above,
-(1.11) $C_{1}(n)=\begin{cases}\frac{2}{3}&n=1,\\ \frac{3\pi}{16}&n=2,\\ \frac{\pi\mathscr{B}(\frac{n}{2}-1,2)}{6\mathscr{B}(\frac{n}{2}-1,5/2)}&n=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n}{2}-1,2)}{16\mathscr{B}(\frac{n}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
+$C_{1}(n)=\begin{cases}\frac{2}{3}&n=1,\\ \frac{3\pi}{16}&n=2,\\ \frac{\pi\mathscr{B}(\frac{n}{2}-1,2)}{6\mathscr{B}(\frac{n}{2}-1,5/2)}&n=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n}{2}-1,2)}{16\mathscr{B}(\frac{n}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$ (1.11)
 Obviously, both quantities $\frac{r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}$ and $\frac{r_{in}(\Omega)}{|I_{min}|}$ are not larger than $1/2$, and therefore can be replaced by $1/2$.
 Moreover, note that if $|\partial(\mathscr{R}\setminus\overline{\Omega})|$ is small compared to $|\mathscr{R}_{n-1}|$, we then have that Pólya’s conjecture holds for the first some eigenvalues.
-Our quantitative estimate is consistent to the asymptotic estimates established by Courant [5] and Courant-Hilbert [6].
 For convex sets, we have a clean and better estimate as following.
@@
 Using Seeley’s asymptotic estimates [37, 38] (see also Hörmander [20]), we first provide an approach to Pólya’s conjecture that might be applied to general smooth domains.
-By a rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$, we mean $\mathscr{R}_{n-1}\subset{\mathbb{R}}^{n-1}$ is a rectangle, and $I\subset{\mathbb{R}}$ is an open interval.
+Definition 1.7 (Strip-tiling domains).
+Let $\Omega$ be a bounded Lipschitz domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$.
+Up to rotations and translations, we assume that $\Omega$ is contained in two hyperplanes $\mathbb{R}^{n-1}\times\{0\}$ and $\mathbb{R}^{n-1}\times\{w(\Omega)\}$ for some $w(\Omega)\in(0,\infty)$.
+We say that $\Omega$ is a strip-tiling domain, if one can use isometries of $\Omega$, obtained by rotating $\Omega$ only in the first n1n-1 directions, to cover $\mathbb{R}^{n-1}\times(0,w(\Omega))$ without overlapped interiors.
+Obviously, a rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I\subset{{{\mathbb{R}}}^{n}}$, $n\geq 2$ is a strip-tiling domain.
+The following Figure 1 provide more examples of general strip-tiling domains.
+Figure 1: Examples of strip-tiling domains in two dimension
+Remark 1.8.
+(i) Note that in the covering we fix the $n$-th direction of $\Omega$, but allow rotating $\Omega$ in the first n1n-1 directions.
+This requires the two side face of $\Omega$ in the $n$-th direction both tile ${\mathbb{R}}^{n-1}$, moreover, the two side face have the same surface measure.
+(ii) Let $\Omega$ be a strip-tiling domain that can cover $\mathbb{R}^{n-1}\times(0,w(\Omega))$.
+Let us denote by $S_{n}(\Omega)$ one of the two side faces of $\Omega$ in the $n$-th direction in what follows.
+Then the measure of $\Omega$ equals $|S_{n}(\Omega)|\times w(\Omega)$.
 Theorem 1.9.
-Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$.
-Assume that $\Omega$ is a piecewise smooth domain contained in $\mathscr{R}$.
-There exists $C_{P}=C(\Omega,n)$ such that if $|\mathscr{R}_{n-1}|\geq C_{P}$ then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\Omega}$.
+Suppose that $\Omega$ is a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$.
+Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction.
+Assume that $\Omega_{0}$ is a piecewise smooth domain contained in $\Omega$.
+There exists $C_{P}=C(\Omega_{0},n)$ such that if $|S_{n}(\Omega)|\geq C_{P}$ then the Pólya conjecture holds on $\Omega\setminus\overline{\Omega}_{0}$.
+Our first observation is that eigenvalues on strip-tiling domains satisfying better estimates than Pólya conjectured.
+Theorem 1.10.
+Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}^{n-1}\times(0,w(\Omega))$, $n\geq 2$.
+Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction.
+Then it holds that for all $k\geq 1$ that
+$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-C_{1}(n-1)|S_{n}(\Omega)|\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},$
+where $C_{1}(n)$ is as in (1.11).
+[…]
+Corollary 1.11.
+[…]
+$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{1}(n-1)}{K}|S_{n}(\tilde{\Omega})|\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}.$
+The above result applies especially to any triangles in the plane.
+Corollary 1.12.
+Let $\Omega$ be a triangle in ${\mathbb{R}}^{2}$ with vertices at $a,b,c$. Then it holds for all $k\geq 1$ that
+$k$ $\leq\frac{|\Omega|}{4\pi}\lambda_{k}(\Omega)-\frac{\max\{|ab|,|bc|,|bc|\}}{12\pi}\lambda_{k}(\Omega)^{\frac{1}{2}}.$
+Remark 1.13.
+Note that previously, it was only known that on an arbitrary triangle $\Omega$,
+$k$ $\leq\frac{|\Omega|}{4\pi}\lambda_{k}(\Omega),\ \forall\ k\geq 1,$
+which is due to Pólya [35]. In special cases of equilateral triangle, see Pinsky [33].
 In what follows, by a cube $Q$ we always mean it is an open cube and denote by $\ell(Q)$ its side length.
@@
 Theorem 1.15.
-Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$.
-Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\mathscr{R}$.
-If $|\mathscr{R}_{n-1}|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
-(1.12) $C_{2}(n)=\begin{cases}2\sqrt{2}\pi,&n=1,\\ 9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
-For $M\geq 2$, by $\tilde{\Omega}$ $M$-tiles $\Omega$ we mean there are $M$ isometries of $\tilde{\Omega}$, $\{\tilde{\Omega}_{k}\}_{1\leq k\leq M}$, with non-overlapped interior, such that $\cup_{1\leq k\leq M}\tilde{\Omega}_{k}\subset\Omega$ and $|\Omega|=M|\tilde{\Omega}|$.
+Let $\Omega$ be a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$.
+Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction.
+Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{\mathbb{R}}}^{n}$}, that is contained in $\Omega$.
+If $|S_{n}(\Omega)|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
+$C_{2}(n)=\begin{cases}2\sqrt{2}\pi,&n=1,\\ 9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$ (1.12)
 Corollary 1.16.
-Let $M\geq 2$.
-Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$, and $\Omega$ $M$-tiles $\mathscr{R}$.
-Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\Omega$.
-If $|\mathscr{R}_{n-1}|\geq C_{2}(n-1)M\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$.
-Note that $M$ isometries of the domain $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$ corresponds to the case that removing $M$ copies of $\{Q_{k}\}_{k\in{\mathbb{N}}}$ from $\mathscr{R}$, the conclusion is obvious.
+Let $\Omega$ be a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$.
+Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction.
+Let $M\geq 2$ and $\Omega_{0}$ $M$-tile $\Omega$.
+Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{\mathbb{R}}}^{n}$}, that is contained in $\Omega_{0}$.
+If $|S_{n}(\Omega)|\geq C_{2}(n-1)M\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $\Omega_{0}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$.
+Note that $M$ isometries of the domain $\Omega_{0}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$ corresponds to the case that removing $M$ copies of $\{Q_{k}\}_{k\in{\mathbb{N}}}$ from $\Omega$, the conclusion is obvious.
 Some further remarks are in order. In what follows for $C>0$ we use $\lfloor C\rfloor$ to denote the integer part of it.
@@
 Then $\mathscr{S}_{Q}\leq 2(2+\sqrt{2})$.
-So for a rectangle $\Omega_{1}\times\Omega_{2}$, $|\Omega_{2}|>1$ and $|\Omega_{1}|>8(1+\sqrt{2})\pi$, one can arbitrarily remove any domain, that is composed by all or some of these cubes $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ with non-overlapped interior, from $\Omega_{1}\times\Omega_{2}$ such that the remaining set satisfies Pólya’s conjecture.
+So for a strip-tiling domain $\Omega$ with $|S_{n}(\Omega)|>8(1+\sqrt{2})\pi$, one can arbitrarily remove any domain, that is composed by all or some of these cubes $\{Q_{1,1},\,\{Q_{k,j}\}_{k\geq 2,1\leq j\leq\ell(k)}\}$ with non-overlapped interior, from $\Omega$ such that the remaining set satisfies Pólya’s conjecture.
 See Figure 2 for some examples.
-(a) Rectangle with a diamond and a room-corridor removed
+(a) A strip-tiling domain with a diamond and a room-corridor removed
 (b) Trapezoid with infinite cubes removed
@@
 1.4 Our method and plan of the paper
-The main strategy to prove Theorem 1.5, Theorem 1.3 and Theorem 1.9 is as following.
+The main strategy of the paper is as following.
 We first use the explicit formula of Riesz means of Dirichlet eigenvalues in $\mathbb{R}$} (cf.
 Lemma 2.1), following the inspiring observation by Laptev [26], to give a refined estimates on product domains $\Omega_{1}\times\Omega_{2}$ (see Theorem 2.3 below).
-This allows us to conclude via monotonicity of Dirichlet eigenvalues (cf.
-[18]) that for any open domains $\Omega$ contained in $\Omega_{1}\times\Omega_{2}$, the first $k_{0}$ Dirichlet eigenvalues satisfies Pólya’s conjecture (cf.
-Corollary 2.4).
-We then establish a quantitative lower bound for the counting function (upper bound for the Dirichlet eigenvalue) on cubes (see Lemma 3.3 and Lemma 3.6 below), which together with our refinements on product domains and monotonicity of Dirichlet eigenvalues to conclude the proof Theorem 1.9.
-In fact, in higher dimensions, $n\geq 3$, the rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I$ can be replaced some more general product $\Omega\times I$, as soon as Pólya’s conjecture holds on $\Omega$, see Theorem 3.7 below.
-Although only a part of discussions in Sections 2 and 3 will be sufficient for the proofs of Theorem 1.3 and Theorem 1.5, the extended discussions are of independent interest and they have other uses.
+Using these refined estimates on product domains together with Pólya’s method [35], we give the refined estimates on strip-tiling domains.
+We then establish a quantitative lower bound for the counting function (upper bound for the Dirichlet eigenvalue) on cubes (see Lemma 3.3 and Lemma 3.6 below), which together with our refinements on strip-tiling domains and monotonicity of Dirichlet eigenvalues to conclude the proof Theorem 1.15.
+Moreover, in higher dimensions, $n\geq 3$, certain product domains can replace the role of strip-tiling domains, see Theorem 3.7 below.
 For Theorem 1.5 we shall apply the Whitney decomposition of a Lipschitz domain $\Omega$, and using the lower bound for the counting function on cubes (see Lemma 3.3 and Lemma 3.6 below) along with an optimization to deduce a lower bound for the counting function $\mathcal{N}_{\Omega}^{D}(\lambda)$ on $\Omega$.
@@
 Finally, let us point it out, our approach depends on a modified min-max argument, which is related to but different from the original ones by Courant [5] and Courant-Hilbert [6].
-By our choose of minimal Admissible rectangles, we do not have to take care of the term in Courant [5] and Courant-Hilbert [6] for cubes which intersect the boundary of $\Omega$.
-The latter is a difficult issue for Courant [5] and Courant-Hilbert [6].
+By our choose of minimal Admissible rectangles, we do not have to use Neumann eigenvalues which is crucial in Courant’s Dirichlet-Neumann bracketing method (see [5, 6, 9, 32]), and gives better constant estimates.
 The paper is organized as follows.
@@
 We shall employ Lemma 3.3 and Lemma 3.6 from Section 3 and Theorem 2.3 from Section 2 to prove the uniform estimates on Lipschitz sets, Theorem 1.5 and its corollary, Theorem 1.3.
-We shall also provide a uniform estimate on general open domains, see Theorem 4.2 below.
 For $C>0$ we use $\lfloor C\rfloor$ to denote the integer part of it.
@@
 Throughout the paper, most of constants are given explicitly, if not we shall point it out.
-Added in remark.
-After the first version of this article was posted on arXiv, we are kindly informed by Frank and Larson, that a similar version to our uniform remainder estimate for Weyl’s law (Theorem 1.5) already appears in [9, Theorem 3.14] and [12, Proposition 3.6], which both show that
-(1.13) $\left|\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|$ $\leq C(n)\Theta_{\Omega}\lambda^{\frac{n-1}{2}}\left(1+\log_{+}(r_{in}(\Omega)\sqrt{\lambda})\right),$
-where $C(n)$ is an implicit constant depending only on the dimension $n$,
-$\Theta_{\Omega}:=\sup_{\delta>0}\delta^{-1}|\left\{x\in\Omega:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)<\delta\right\}|<\infty,$
-$\Theta_{\Omega}$ plays the same role of $C_{Lip}(\Omega)|\partial\Omega|$ in our estimate.
-In particular, it appears from the proof of [9, Theorem 3.14] that the above constant $C(n)$ can also be calculated.
-The proof of [12, Proposition 3.6] uses Tauberian theory, while the proof of [9, Theorem 3.14] uses Courant’s method of Dirichlet-Neumann bracketing and some refined estimates for cubes intersecting the boundary.
-Our proof for estimates of the reminder of Weyl’s law (Theorem 1.5) uses the minimal Admissible rectangle (see Theorem 2.3 below for refined eigenvalue estimates on product domains), a subtracting argument that allows us to avoid the cubes intersecting the boundary, and then lower bound for the counting function of Dirichlet eigenvalues.
-Moreover, our method has an advantage that gives many “irregular shaped domains” that verify the Pólya conjecture (see Theorem 1.9, Remark 1.11 and Theorem 3.7 below).
-2 Refined eigenvalue estimates on product domains
-In this section, we prove refined eigenvalue estimates on product domains, Theorem 2.3 and its corollary, Corollary 2.4.
-Recall that, if an open domain $\Omega$ is in the real line, then its eigenvalues have explicit lower bounds, which satisfy for $k\geq 1$
+2 Refined eigenvalue estimates
+In this section, we prove refined eigenvalue estimates on product domains and strip-tiling domains.
+2.1 Product domains
+We first prove Theorem 2.3 and its corollary, Corollary 2.4. Recall that, if an open domain $\Omega$ is in the real line, then its eigenvalues have explicit lower bounds, which satisfy for $k\geq 1$
 $\lambda_{k}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}}k^{2}.$
@@
 Let $\Omega\subset{\mathbb{R}}$ be bounded. It holds that
-(2.1) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $\leq\frac{2|\Omega|}{3\pi}\lambda^{3/2}\min\left\{\left(1-\frac{3\pi}{16|\Omega|\sqrt{\lambda}}\right)_{+},\,\left(1-\left(\frac{\lfloor\frac{|\Omega|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega|\sqrt{\lambda}}\right)\right\},$
+$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $\leq\frac{2|\Omega|}{3\pi}\lambda^{3/2}\min\left\{\left(1-\frac{3\pi}{16|\Omega|\sqrt{\lambda}}\right)_{+},\,\left(1-\left(\frac{\lfloor\frac{|\Omega|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega|\sqrt{\lambda}}\right)\right\},$ (2.1)
 and for $\lambda>\frac{\pi^{2}}{|\Omega|^{2}}$,
-(2.2) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\leq\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\leq\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
+$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\leq\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\leq\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$ (2.2)
 Moreover, if $\Omega$ is a connected open interval, it holds for $\lambda>\lambda_{1}(\Omega)=\frac{\pi^{2}}{|\Omega|^{2}}$ that
-(2.3) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\geq\frac{|\Omega|}{4}\lambda\left(1-\frac{2\pi}{|\Omega|\sqrt{\lambda}}+\frac{\pi^{2}}{|\Omega|^{2}\lambda}\right).$
+$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\geq\frac{|\Omega|}{4}\lambda\left(1-\frac{2\pi}{|\Omega|\sqrt{\lambda}}+\frac{\pi^{2}}{|\Omega|^{2}\lambda}\right).$ (2.3)
 Proof.
@@
 Let us consider the ball $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$.
-For each $1\leq k\leq k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$, $(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}$ equals the area of the rectangle $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, which is contained in $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$.
+For each $1\leq k\leq k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$, $(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}$ equals the area of the rectangle $k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, which is contained in $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$.
 Moreover, for each $k$, there is at least a triangle with vertices at $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-(k-1)^{2})^{1/2})$, $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ and $(k,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ contained in the ball.
 Moreover, if $k_{\lambda}\neq\frac{|\Omega|}{\pi}\sqrt{\lambda}$, there is another triangle insider the 1/4 ball with vertices at $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$, $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$ and $(\frac{|\Omega|}{\pi}\sqrt{\lambda},0)$.
-See Figure $2$Picture drawn by Xinyi Jiang..
+See Figure 3222Picture drawn by Xinyi Jiang..
 Figure 3: The 1/2 sum is controlled by the 1/4 ball
@@
 $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $=\sum_{1\leq k\leq k_{\lambda}}(\lambda-\frac{\pi^{2}}{|\Omega|^{2}}k^{2})^{1/2}=\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}.$
-Note that the union of the rectangles $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, $1\leq k\leq k_{\lambda}$, contains the ball $B(0,\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda}+1})$ intersecting the first quadrant, since
+Note that the union of the rectangles $k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, $1\leq k\leq k_{\lambda}$, contains the ball $B(0,\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda}+1})$ intersecting the first quadrant, since
 $\sqrt{(k-1)^{2}+(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})}$ $=\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2k+1}\geq\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2k_{\lambda}+1}$
@@
 where
-(2.4) $C_{3}(p)=\begin{cases}\frac{1}{8},&p=1,\\ \frac{1}{9},&1<p\leq 2,\\ \frac{1}{8}\left(\frac{247}{256}\right)^{p-1},&p>2.\end{cases}$
+$C_{3}(p)=\begin{cases}\frac{1}{8},&p=1,\\ \frac{1}{9},&1<p\leq 2,\\ \frac{1}{8}\left(\frac{247}{256}\right)^{p-1},&p>2.\end{cases}$ (2.4)
 Proof.
 By the semigroup type property of the Riesz means and (2.1), we have
-[…]
-[…]
-[…]
+$\sum_{k:\,\lambda_{k}<\lambda}(\lambda-\lambda_{k})^{p}$ $\leq\mathscr{B}(p-1,2)^{-1}\sum_{k}\int_{0}^{\infty}v^{p-2}(\lambda-\lambda_{k}-v)_{+}\,dv$
+$\leq\frac{2|\Omega|}{3\pi\mathscr{B}(p-1,2)}\int_{0}^{\infty}v^{p-2}(\lambda-v)_{+}^{3/2}\left(1-\frac{3\pi}{16|\Omega|\sqrt{(\lambda-v)_{+}}}\right)_{+}\,dv$
+$\leq\frac{2|\Omega|}{3\pi\mathscr{B}(p-1,2)}\lambda^{p+\frac{1}{2}}\mathscr{B}(p-1,5/2)-\frac{1}{8\mathscr{B}(p-1,2)}\int_{0}^{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}v^{p-2}(\lambda-v)_{+}\,dv.$ (2.5)
 Noticing that $\lambda_{1}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}},$ so for $\lambda>\lambda_{1}(\Omega)$,
@@
 For $1<p\leq 2$,
-[…]
+$\frac{1}{8\mathscr{B}(p-1,2)}\int_{0}^{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}v^{p-2}(\lambda-v)_{+}\,dv\geq\frac{\lambda^{p}}{8\mathscr{B}(p-1,2)}\frac{249}{256}\mathscr{B}(p-1,2)\geq\frac{\lambda^{p}}{9}.$ (2.6)
 For $p>2$, we have
-[…]
+$\int_{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}^{\lambda}v^{p-2}(\lambda-v)_{+}\,dv$ $=\frac{\lambda^{p}}{p-1}-\frac{\lambda^{p}}{p}-\left(\frac{\lambda(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}})^{p-1}}{p-1}-\frac{(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}})^{p}}{p}\right)$
 $=\frac{\lambda^{p}}{p(p-1)}-\left(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}\right)^{p-1}\left(\frac{\lambda}{p-1}-\frac{(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}})}{p}\right)$
@@
 and therefore
-[…]
-[…]
+$\frac{1}{8\mathscr{B}(p-1,2)}\int_{0}^{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}v^{p-2}(\lambda-v)_{+}\,dv$
+$=\frac{1}{8\mathscr{B}(p-1,2)}\left(\int_{0}^{\lambda}v^{p-2}(\lambda-v)_{+}\,dv-\int_{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}^{\lambda}v^{p-2}(\lambda-v)_{+}\,dv\right)$
 $\geq\frac{1}{8\mathscr{B}(p-1,2)}\left(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}\right)^{p-1}\frac{\lambda}{p(p-1)}$
@@
 where $C_{1}(n_{1})$ is as in (1.11),
-(2.7) $C_{1}(n_{1})=\begin{cases}\frac{2}{3}&n_{1}=1,\\ \frac{3\pi}{16}&n_{1}=2,\\ \frac{\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{6\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}&n_{1}=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{16\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n_{1}-2}{2}}&n_{1}\geq 5.\end{cases}$
+$C_{1}(n_{1})=\begin{cases}\frac{2}{3}&n_{1}=1,\\ \frac{3\pi}{16}&n_{1}=2,\\ \frac{\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{6\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}&n_{1}=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{16\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n_{1}-2}{2}}&n_{1}\geq 5.\end{cases}$ (2.7)
 Proof.
@@
 Then for all $\lambda>\lambda_{1}(\Omega)$, where
-(2.8) $\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{1})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{1}|^{2/n}\omega(n_{1})^{2/n_{1}}}+\frac{\pi^{2}}{|\Omega_{2}|^{2}},$
+$\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{1})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{1}|^{2/n}\omega(n_{1})^{2/n_{1}}}+\frac{\pi^{2}}{|\Omega_{2}|^{2}},$ (2.8)
 we have
@@
 $\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n_{1})}{|\Omega_{2}|\sqrt{\lambda}}\right)$
-(2.9) $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$
+$\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$ (2.9)
 Recall that
 $k_{0}=\left\lfloor\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n_{1})}{|\Omega_{2}|}\right]^{n}\right\rfloor\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{0}^{n/2}.$
-Let $0\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
+Let $0\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2.1) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
 $k\leq\mathcal{N}^{D}_{\Omega}(\lambda_{k}(\Omega)+\epsilon)$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}(\lambda_{k}(\Omega)+\epsilon)^{n/2}.$
@@
 34) Dirichlet eigenvalues satisfy Pólya’s conjecture.
+2.2 Strip-tiling domains
+We next use Theorem 2.3 together with Pólya’s original method to prove Theorem 1.10.
+Proof of Theorem 1.10.
+Given a strip-tiling domain $\Omega$, let us consider $\ell^{n-1}$ isometries of $\Omega$, $\{\Omega_{j}\}_{1\leq j\leq\ell^{n-1}}$ that are placed as the tiling procedure such that in the first n1n-1 directions, each line only intersects at most $\ell$ isometries of $\Omega$.
+Consider the minimal Admissible rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times(0,w(\Omega))$ that contains these $\ell^{n-1}$ isometries of $\Omega$.
+Then we have
+(i) $|\mathscr{R}_{n-1}|\geq\ell^{n-1}|S_{n}(\Omega)|$,
+(ii) $|\mathscr{R}\setminus(\cup_{j=1}^{\ell^{n-1}}\overline{\Omega_{j}})|\leq C(\Omega)\ell^{n-2}$.
+This is true since $\Omega$ tiles the strip ${\mathbb{R}}^{n-1}\times[0,w(\Omega)]$, and the set $\mathscr{R}\setminus(\cup_{j=1}^{\ell^{n-1}}\overline{\Omega_{j}})$ locates near the 2(n1)2(n-1) faces of $\mathscr{R}$ in the first n1n-1 directions.
+See the following Figure 4.
+Figure 4: Examples of strip-tiling domains in two dimension
+By Theorem 2.3, we see that for each $k\geq 1$,
+$\ell^{n-1}N^{D}_{{\Omega}}(\lambda)$ $\leq N^{D}_{\cup_{j=1}^{\ell^{n-1}}\overline{\Omega_{j}}}(\lambda)\leq N^{D}_{\mathscr{R}}(\lambda)$
+$\leq\frac{|\mathscr{R}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|w(\Omega)|\sqrt{\lambda}}\right)_{+},$
+which implies that
+$N^{D}_{{\Omega}}(\lambda)$ $\leq\frac{\omega(n)}{(2\pi)^{n}}\frac{|\mathscr{R}|}{\ell^{n-1}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|w(\Omega)|\sqrt{\lambda}}\right)_{+}.$
+Sending $\ell$ to $\infty$, using the properties (i) and (ii) of $\mathscr{R}$, we see that for each $k\geq 1$,
+$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-C_{1}(n-1)|S_{n}(\Omega)|\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}.$
+The proof is complete. ∎
+Remark 2.5.
+In Theorem 2.3 and Theorem 1.10, we only keep one lower order term.
+One can have a better estimate especially for large eigenvalues, by keeping more lower order term.
+We present it below in two and three dimensions.
+Let $\Omega_{1}\subset{\mathbb{R}}^{n_{1}}$ and $\Omega_{2}\subset{\mathbb{R}}$ be bounded open domains, $n_{1}=1,2$. If $n_{1}=1$, then by (2.2), we have
+$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|}{\pi}(\lambda-\lambda_{j}(\Omega_{2}))^{1/2}\leq\frac{|\Omega_{1}|}{\pi}\left(\frac{|\Omega_{2}|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega_{2}|}\sqrt{\lfloor\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+}$
+$\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\lfloor\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+}.$
+If $n_{1}=2$, we assume that Pólya’s conjecture holds on $\Omega_{1}$. By (2.1) we have
+$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|\omega(2)}{(2\pi)^{2}}(\lambda-\lambda_{j}(\Omega_{2}))$
+$\leq\frac{|\Omega_{1}||\Omega_{2}|}{6\pi^{2}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
+Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}^{n-1}\times(0,w(\Omega))$, $n=2,3$, with $S_{n}(\Omega)$ be a face in the $n$-direction.
+From the proof of Theorem 1.10 we deduce similar estimate as following.
+For the case $n=2$, it holds that
+$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{2}{|w(\Omega)|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|w(\Omega)|^{2}\lambda}\sqrt{\lfloor\frac{|w(\Omega)|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+},$
+and for $n=3$, it holds that
+$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{6\pi^{2}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|w(\Omega)|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|w(\Omega)|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|w(\Omega)|\sqrt{\lambda}}\right)_{+}.$
+Similar to Corollary 2.4 we have the following perturbation result.
+Corollary 2.6.
+Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}^{n-1}\times(0,w(\Omega))$, $n\geq 2$.
+Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction.
+Let $\Omega_{0}$ be an open set contained in $\Omega$, and
+$k_{0}=\left\lfloor\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega_{0}|}{|\Omega|}\right)^{-1}\frac{C_{1}(n-1)}{|w(\Omega)|}\right]^{n}\right\rfloor.$
+Then for each $0\leq k\leq k_{0}$, the Pólya conjecture holds on $\Omega$ for $\lambda_{k}(\Omega)$, i.e.,
+$k\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$
 3 Pólya’s conjecture on a class of domains
@@
 Theorem 3.1.
-Suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture, $n\geq 2$, $\Omega_{2}$ is an interval in ${\mathbb{R}}$ and $\Omega_{3}$ is a piecewise smooth domain contained in $\Omega_{1}\times\Omega_{2}$.
+(i) Suppose that $\Omega$ is a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$.
+Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction.
+Suppose that $\Omega_{3}$ is a piecewise smooth domain contained in $\Omega$.
+There exists $C_{P}=C(\Omega_{3},n)$ such that if $|S_{n}(\Omega)|\geq C_{P}$ then the Pólya conjecture holds on $\Omega\setminus\overline{\Omega_{3}}$.
+(ii) Suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture, $n\geq 2$, $\Omega_{2}$ is an interval in $\mathbb{R}$} and $\Omega_{3}$ is a piecewise smooth domain contained in $\Omega_{1}\times\Omega_{2}$.
 There exists $C_{P}=C(\Omega_{3},n)$ such that if $|\Omega_{1}|\geq C_{P}$ then the Pólya conjecture holds on $\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$.
 Proof.
+We only prove (ii), the proof of (i) is the same by using Theorem 1.10.
 Let $\Omega=\Omega_{1}\times\Omega_{2}\setminus\overline{\Omega_{3}}$.
@@
 It follows from Seeley [37, 38] (see also [20, Corollary 17.5.11 & (17.5.16)’]) that for some $\Lambda>0$, it holds for $\lambda>\Lambda$ that
-(3.1) $\left|\mathcal{N}^{D}_{\Omega_{3}}(\lambda)-\frac{|\Omega_{3}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|\leq C(\Omega_{3})\lambda^{\frac{n-1}{2}}.$
+$\left|\mathcal{N}^{D}_{\Omega_{3}}(\lambda)-\frac{|\Omega_{3}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|\leq C(\Omega_{3})\lambda^{\frac{n-1}{2}}.$ (3.1)
 We assume that
@@
 Note that (3.1) can be stated for all $\lambda>0$ without referring to $\Lambda$, but generally with larger constant $C(\Omega_{3})$.
-We state in this way to show that the combined argument, that is for small eigenvalues using Corollary 2.4 while for large constant using the subtracting argument, will give better estimate, see Remark 3.5 below.
+We state in this way to show that the combined argument, that is for small eigenvalues using Corollary 2.4 while for large eigenvalues using the subtracting argument, will give better estimate, see Remark 3.5 below.
 We next prove the concrete examples of domains obtained by removing cubes, Theorem 1.15. Let us start with the two dimensional case.
@@
 We have
-$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}),\,a,b\in\mathbb{N}\right\}.$
+$\mathcal{N}_{Q}^{D}(\lambda)$ $=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}),\,a,b\in\mathbb{N}\right\}.$
 So the number $\mathcal{N}_{Q}^{D}(\lambda)$ is not less than the number of the unit cubes contained in the first quadrant intersecting the ball $B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{2})$, where $\sqrt{2}$ is the diameter of the unit cube.
@@
 Remark 3.4.
-Using (2.3) one can give a lower bound for $\mathcal{N}^{D}_{\mathcal{R}}(\lambda)$ for any rectangle $\mathcal{R}$ in ${\mathbb{R}}^{2}$.
+Using (2.3) one can give a lower bound for $\mathcal{N}^{D}_{\mathscr{R}}(\lambda)$ for any rectangle $\mathscr{R}$ in ${\mathbb{R}}^{2}$.
 Inductively, a lower bound for the eigenvalue counting function can be explicitly given on any cuboid in higher dimension.
@@
 Proof of the case $n=2$ in Theorem 1.15.
-Denote by $\Omega_{0}=\Omega_{1}\times\Omega_{2}$ and $\Omega=\Omega_{1}\times\Omega_{2}\setminus\overline{\cup_{k=1}^{\infty}Q_{k}}$. Let
+Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}\times(0,w(\Omega))$, and let $S_{n}(\Omega)$ be the surface of a side face of $\Omega$ in the $n$-th direction.
+Denote by $\Omega_{0}=\Omega\setminus\overline{\cup_{k=1}^{\infty}Q_{k}}$.
+Let
 $\mathscr{V}_{Q}:=\left|\cup_{k=1}^{\infty}Q_{k}\right|=\sum_{k=1}^{\infty}\ell(Q_{k})^{n}.$
-Then
+Then by the definition of $\{Q_{k}\}$, it holds that $w(\Omega)\geq 1$ and
 $\mathscr{V}_{Q}\leq\mathscr{S}_{Q}<\infty.$
 By Theorem 1.10 we have for $\lambda>0$ that
-$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{{2}}{3|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
+$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{{2}}{3|w(\Omega)|\sqrt{\lambda}}\right)_{+}.$
 by Corollary 2.6 we have for
-$k_{0}=\left\lfloor\frac{|\Omega|}{4\pi}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{2}{3|\Omega_{2}|}\right]^{2}\right\rfloor=\left\lfloor\frac{|\Omega|}{4\pi}\left[\frac{2|\Omega_{1}|}{3\mathscr{V}_{Q}}\right]^{2}\right\rfloor,$
-the first $k_{0}$ Dirichlet eigenvalues of $\Omega$ satisfy Pólya’s conjecture, that is, for $1\leq i\leq k_{0}$,
-$\frac{|\Omega|\omega(2)}{(2\pi)^{2}}\lambda_{i}(\Omega)\geq i.$
-In particular, since $|\Omega_{1}|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}\geq 2\sqrt{2}\pi\mathscr{V}_{Q}$,
-$|\Omega|=|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1)\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1),$
+$k_{0}=\left\lfloor\frac{|\Omega_{0}|}{4\pi}\left[\left(1-\frac{|\Omega_{0}|}{|\Omega|}\right)^{-1}\frac{2}{3|w(\Omega)|}\right]^{2}\right\rfloor=\left\lfloor\frac{|\Omega_{0}|}{4\pi}\left[\frac{2|S_{n}(\Omega)|}{3\mathscr{V}_{Q}}\right]^{2}\right\rfloor,$
+the first $k_{0}$ Dirichlet eigenvalues of $\Omega_{0}$ satisfy Pólya’s conjecture, that is, for $1\leq i\leq k_{0}$,
+$\frac{|\Omega_{0}|\omega(2)}{(2\pi)^{2}}\lambda_{i}(\Omega)\geq i.$
+In particular, since $|S_{n}(\Omega)|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}\geq 2\sqrt{2}\pi\mathscr{V}_{Q}$,
+$|\Omega_{0}|=|\Omega|-\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1)\mathscr{V}_{Q}\geq(2\sqrt{2}\pi-1),$
 we find that
-$\lambda_{k_{0}}(\Omega)\geq\frac{4\pi}{|\Omega|}\left(\frac{|\Omega|}{4\pi}\left[\frac{2|\Omega_{1}|}{3\mathscr{V}_{Q}}\right]^{2}-1\right)>33.$
-Recall that Lemma 2.1 implies that for $\lambda>\pi^{2}/|\Omega_{1}|^{2}+\pi^{2}/|\Omega_{2}|^{2}$,
-$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|}{\pi}(\lambda-\lambda_{j}(\Omega_{2}))^{1/2}$
-$\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\lfloor\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right).$
-In particular, for $\lambda>33$, it holds that
-$\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\frac{3}{2}\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}}\right)$
-(3.2) $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{1}{|\Omega_{2}|\sqrt{\lambda}}\right).$
+$\lambda_{k_{0}}(\Omega)\geq\frac{4\pi}{|\Omega|}\left(\frac{|\Omega|}{4\pi}\left[\frac{2|S_{n}(\Omega)|}{3\mathscr{V}_{Q}}\right]^{2}-1\right)>33.$
+By Remark 2.5, we have
+$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{2}{|w(\Omega)|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|w(\Omega)|^{2}\lambda}\sqrt{\lfloor\frac{|w(\Omega)|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+}.$
+This together with $|w(\Omega)|\geq 1$ implies that for $\lambda>33$, it holds
+$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{1}{|w(\Omega)|\sqrt{\lambda}}\right).$ (3.2)
 Now consider the union of open sets $\Omega$, $Q_{k}$, $k\in{\mathbb{N}}$, which are disjoint and contained in $\Omega_{1}\times\Omega_{2}$, but satisfy
-$|\Omega_{1}\times\Omega_{2}|=|\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}|.$
-So by monotonicity of Dirichlet eigenvalues, Pólya’s conjecture holds on $\Omega\cup\cup_{1\leq k<\infty}Q_{k}$,
-$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{0}}(\lambda).$
-Moreover, by Lemma 3.3 (ii), we see that for $\lambda\geq\lambda_{k_{0}}(\Omega)$,
-$\mathcal{N}^{D}_{\Omega\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{k}}(\lambda)$
-$\geq\mathcal{N}^{D}_{\Omega}(\lambda)+\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right).$
+$|\Omega|=|\Omega_{0}\cup\cup_{k\in{\mathbb{N}}}Q_{k}|.$
+So by monotonicity of Dirichlet eigenvalues, Pólya’s conjecture holds on $\Omega_{0}\cup\cup_{1\leq k<\infty}Q_{k}$, and
+$\mathcal{N}^{D}_{\Omega_{0}\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)\leq\mathcal{N}^{D}_{\Omega}(\lambda).$
+Moreover, by Lemma 3.3 (ii), we see that for $\lambda>0$,
+$\mathcal{N}^{D}_{\Omega_{0}\cup\cup_{k\in{\mathbb{N}}}Q_{k}}(\lambda)$ $=\mathcal{N}^{D}_{\Omega_{0}}(\lambda)+\sum_{k\in{\mathbb{N}}}\mathcal{N}^{D}_{Q_{k}}(\lambda)$
+$\geq\mathcal{N}^{D}_{\Omega_{0}}(\lambda)+\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right).$
 Combining the above two estimates, we deduce that for $\lambda\geq\lambda_{k_{0}}(\Omega)>33$,
-$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{{1}}{|\Omega_{2}|\sqrt{\lambda}}\right)-\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right)$
-$=\frac{|\Omega_{1}||\Omega_{2}|-\mathscr{V}_{Q}}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|\Omega_{1}|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right)$
-$=\frac{|\Omega|}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|\Omega_{1}|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right).$
-Since $|\Omega_{1}|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}$, we see that
-$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda,$
+$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{1}{|w(\Omega)|\sqrt{\lambda}}\right)-\sum_{k\in{\mathbb{N}}}\left(\frac{\ell(Q_{k})^{2}}{4\pi}\lambda-\frac{\sqrt{2}\ell(Q_{k})}{2}\sqrt{\lambda}\right)$
+$=\frac{|\Omega|-\mathscr{V}_{Q}}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|S_{n}(\Omega)|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right)$
+$=\frac{|\Omega_{0}|}{4\pi}\lambda-\sqrt{\lambda}\left(\frac{|S_{n}(\Omega)|}{4\pi}-\frac{\sqrt{2}}{2}\mathscr{S}_{Q}\right).$
+Since $|S_{n}(\Omega)|\geq 2\sqrt{2}\pi\mathscr{S}_{Q}$, we see that
+$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)$ $\leq\frac{|\Omega_{0}|}{4\pi}\lambda,$
 which completes the proof. ∎
@@
 (i) For $n=3$, we have
-$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}+c^{2}),\,a,b,c\in\mathbb{N}\right\}.$
+$\mathcal{N}_{Q}^{D}(\lambda)$ $=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}(a^{2}+b^{2}+c^{2}),\,a,b,c\in\mathbb{N}\right\}.$
 So the number $\mathcal{N}_{Q}^{D}(\lambda)$ is not less than the number of the unit cubes contained in the first quadrant intersecting the ball $B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{3})$, where $\sqrt{3}$ is the diameter of the unit cube.
@@
 (ii) For $n\geq 4$, we also have
-$\mathcal{N}_{Q}^{D}(\lambda)$ =#$=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}\sum_{i=1}^{n}a_{i}^{2},\,a_{i}\in\mathbb{N}\right\}.$
+$\mathcal{N}_{Q}^{D}(\lambda)$ $=\#\left\{\lambda_{k}<\lambda:\,\lambda_{k}=\frac{\pi^{2}}{\ell(Q)^{2}}\sum_{i=1}^{n}a_{i}^{2},\,a_{i}\in\mathbb{N}\right\}.$
 So the number $\mathcal{N}_{Q}^{D}(\lambda)$ is not less than the number of the unit cubes contained in the first quadrant intersecting the ball $B(0,\ell(Q)\sqrt{\lambda}/\pi-\sqrt{n})$, where $\sqrt{n}$ is the diameter of the unit cube.
@@
 Recall that by (1.11) and (2.4), it holds for $n\geq 4$ that
-(3.3) $C_{1}(n-1)=\frac{3\pi\mathscr{B}(\frac{n-1}{2}-1,2)}{2\mathscr{B}(\frac{n-1}{2}-1,5/2)}C_{3}(\frac{n-1}{2})=\frac{3\pi\mathscr{B}(\frac{n-3}{2},2)}{2\mathscr{B}(\frac{n-3}{2},5/2)}C_{3}(\frac{n-1}{2}),$
+$C_{1}(n-1)=\frac{3\pi\mathscr{B}(\frac{n-1}{2}-1,2)}{2\mathscr{B}(\frac{n-1}{2}-1,5/2)}C_{3}(\frac{n-1}{2})=\frac{3\pi\mathscr{B}(\frac{n-3}{2},2)}{2\mathscr{B}(\frac{n-3}{2},5/2)}C_{3}(\frac{n-1}{2}),$ (3.3)
 where
@@
 Theorem 3.7.
-Let $n\geq 3$ and suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture (e.g.
+Let $n\geq 3$.
+(i) Suppose that $\Omega_{1}\subset{\mathbb{R}}^{n-1}$ satisfies Pólya’s conjecture (e.g.
 ball, tiling domain, or our example in ${\mathbb{R}}^{n-1}$), $\Omega_{2}$ is an interval in $\mathbb{R}$}.
@@
 $C_{2}(n)=\begin{cases}9\sqrt{3},&n=2,\\ \frac{6(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{\mathscr{B}(\frac{n}{2}-1,2)}&n=3,4,\\ \frac{16(n+1)^{3/2}\mathscr{B}(\frac{n}{2}-1,5/2)}{3\mathscr{B}(\frac{n}{2}-1,2)}\left(\frac{256}{247}\right)^{\frac{n-2}{2}}&n\geq 5.\end{cases}$
+(ii) Let $\Omega$ be a strip-tiling domain in ${{\mathbb{R}}}^{n}$}, $n\geq 2$.
+Let $S_{n}(\Omega)$ be the surface of a side face of $\Omega$ in the $n$-th direction.
+Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{\mathbb{R}}}^{n}$} that is contained in $\Omega$.
+If $|S_{n}(\Omega)|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then Pólya’s conjecture holds on $\Omega\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$,
 Proof.
+We only prove (i), the proof of (ii) is the same by using Theorem 1.10.
 Denote by $\Omega_{0}=\Omega_{1}\times\Omega_{2}$, $\Omega_{3}$ the union $\cup_{k\in{\mathbb{N}}}Q_{k}$}, and $\Omega=\Omega_{0}\setminus\overline{\Omega_{3}}$. Let
@@
 $\lambda_{k_{0}}(\Omega)\geq\left(6\pi^{2}\frac{k_{0}}{|\Omega|}\right)^{2/3}>\left(6\pi^{2}\frac{\frac{9\pi|\Omega_{1}|^{3}|\Omega|}{2^{13}\mathscr{V}_{Q}^{3}}-1}{|\Omega|}\right)^{2/3}>\left(6\pi^{2}\frac{\frac{8\pi|\Omega_{1}|^{3}|\Omega|}{2^{13}\mathscr{V}_{Q}^{3}}}{|\Omega|}\right)^{2/3}>61>3\pi^{2}.$
-Recall that Lemma 2.1 implies that for $\lambda>\lambda_{1}(\Omega_{2})$,
+Recall that Remark 2.5 implies that for $\lambda>\lambda_{1}(\Omega_{2})$,
 $\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|\omega(2)}{(2\pi)^{2}}(\lambda-\lambda_{j}(\Omega_{2}))$
-(3.4) $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right).$
+$\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right).$ (3.4)
 Since $\lambda_{k}(\Omega)>61$, $|\Omega_{2}|\geq 1$, we have that
@@
 and hence,
-(3.5) $\mathcal{N}^{D}_{\Omega_{0}}(\lambda)=\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{\pi}{3|\Omega_{2}|\sqrt{\lambda}}\right).$
+$\mathcal{N}^{D}_{\Omega_{0}}(\lambda)=\mathcal{N}^{D}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $\leq\frac{|\Omega_{1}||\Omega_{2}|\omega(3)}{(2\pi)^{3}}\lambda^{3/2}\left(1-\frac{\pi}{3|\Omega_{2}|\sqrt{\lambda}}\right).$ (3.5)
 For
@@
 $k_{0}$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega|}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n-1)}{|\Omega_{2}|}\right]^{n}=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\left[\left(\frac{\mathscr{V}_{Q}}{|\Omega_{1}||\Omega_{2}|}\right)^{-1}\frac{C_{1}(n-1)}{|\Omega_{2}|}\right]^{n}$
-(3.6) $=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\frac{(C_{1}(n-1)|\Omega_{1}|)^{n}}{\mathscr{V}_{Q}^{n}},$
+$=\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\frac{(C_{1}(n-1)|\Omega_{1}|)^{n}}{\mathscr{V}_{Q}^{n}},$ (3.6)
 the Pólya conjecture holds true for $1\leq k\leq k_{0}$ on $\Omega$, i.e.,
@@
 Then it holds on $\Omega=\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}}$ that
-(3.7) $k\leq\begin{cases}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},\,&\,k\leq 2,\\ \frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},\,&\,k>2.\end{cases}$
+$k\leq\begin{cases}\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2},\,&\,k\leq 2,\\ \frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}},\,&\,k>2.\end{cases}$ (3.7)
 Proof.
@@
 This is equivalent to that, for $k\geq 3$,
-(3.8) $k\leq\frac{|\Omega|}{(2\pi)^{n}}\omega(n)\lambda_{k}^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}^{\frac{n-1}{2}},$
+$k\leq\frac{|\Omega|}{(2\pi)^{n}}\omega(n)\lambda_{k}^{n/2}+\frac{\omega(n)n^{3/2}\pi\mathscr{S}_{Q}}{(2\pi)^{n}}\lambda_{k}^{\frac{n-1}{2}},$ (3.8)
 which completes the proof. ∎
@@
 The proof is similar to the proof of Theorem 2.3. Let $\Omega_{0}=\Omega_{1}\setminus\overline{\cup_{j\in{\mathbb{N}}}Q_{j}}$. Then for all $\lambda>\lambda_{1}(\Omega)$, where
-(3.9) $\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{0})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{0}|^{2/n}\omega(n)^{2/n}}+\frac{\pi}{|\Omega_{2}|^{2}},$
+$\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{0})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{0}|^{2/n}\omega(n)^{2/n}}+\frac{\pi}{|\Omega_{2}|^{2}},$ (3.9)
 we have via Theorem 3.8 that
@@
 which completes the proof. ∎
-4 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
+4 Quantitative remainder estimate
 In this section, we prove our main uniform estimates on Lipschitz domains. Let us recall the Whitney decomposition, see [16, Appendix J].
@@
 We therefore have
-(4.1) $\mathcal{N}_{\Omega}^{D}(\lambda)$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\left(\epsilon\lambda^{1/2}+5n^{2}\pi\log_{2}\frac{10r_{in}}{\epsilon}\right).$
+$\mathcal{N}_{\Omega}^{D}(\lambda)$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\left(\epsilon\lambda^{1/2}+5n^{2}\pi\log_{2}\frac{10r_{in}}{\epsilon}\right).$ (4.1)
 Let $\mathscr{R}=\mathscr{R}_{n-1}\times I_{min}$ be the minimal rectangle for $\Omega$. Since
@@
 $\left|\{x\in\Omega_{1}:\,\mathrm{dist}(x,\partial\Omega_{1})<\epsilon\}\right|$ $\leq\epsilon|\partial\mathscr{R}\cap\partial\Omega_{1}|+C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}\cap(\partial\Omega\setminus\partial\mathscr{R})|$
-(4.2) $\leq C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}|.$
+$\leq C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}|.$ (4.2)
 Recall also that $r_{in}(\Omega_{1})$ denotes the inner radius of $\Omega_{1}$.
@@
 $\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda}}{\pi}\geq\frac{10}{\pi}>e,$
-and therefore $f(\Lambda)$ is monotone decreasing on $[\frac{1}{\mathrm{width}(\Omega)^{2}},\infty)$.
+and therefore $f(\Lambda)$ is monotone decreasing on $\frac{1}{\mathrm{width}(\Omega)^{2}},\infty)$.
 Note that for $\Lambda=\frac{1}{\mathrm{width}(\Omega)^{2}}$,
@@
 The remaining proof is the same. The proof is complete. ∎
-For open sets, we also have a uniform estimate.
-Let us remark that, since the boundary of an open set may be complicate, in general one can not have a uniform estimate in term of its boundary as in Theorem 1.5; see [9, Theorem 3.14] for similar but earlier result.
-Note that $C_{1}(n-1)$ is defined via (1.11).
-Theorem 4.2.
-Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded open domain, $n\geq 2$, let $Q$ be a cube that contains $\Omega$. For any $20<\epsilon<\ell(Q)/$ 2, it holds for all $k\in{\mathbb{N}}$ that
-(4.3) $k\leq\frac{|\Omega_{\epsilon}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+\frac{\omega(n)\lambda^{\frac{n-1}{2}}}{(2\pi)^{n}}\left(2n^{2}\pi\frac{|Q\setminus\Omega_{\epsilon}|}{\epsilon}-C_{1}(n-1)\ell(Q)^{n-1}\right).$
-where $\Omega_{\epsilon}=\{x\in Q:\,\mathrm{dist}(x,\Omega)<\epsilon\}$. In particular, one can take $Q$ to be of side length $\mathrm{diam}(\Omega)$.
-Proof.
-Since $\epsilon<\ell(Q)/2$, we can divide $\mathscr{R}$ into the union of cubes $Q_{j}$, which are cubes of side length $2^{-k}\ell(Q)$, where $k\in\mathbb{Z}_{+}$ is such that
-(4.4) $\epsilon/2<2^{-k}\ell(Q)\sqrt{n}\leq\epsilon.$
-Denote by $\tilde{Q}$ the union of all $Q_{j}$, such that $Q_{j}$ entirely locates in $Q\setminus\overline{\Omega}$. Then the number of $Q_{j}$ that composes $\tilde{Q}$ is controlled by
-#$\#\{Q_{j}:\,Q_{j}\subset\tilde{Q}\}\leq\frac{|Q\setminus\Omega|}{2^{-kn}\ell(Q)^{n}}.$
-On the other hand, $\Omega\subset\mathscr{R}\setminus\tilde{Q}$, which is further contained in
-$\{x\in Q:\,\mathrm{dist}(x,\Omega)<2^{-k}\ell(Q)\sqrt{n}\}\subset\Omega_{\epsilon}.$
-By Theorem 2.3, we have
-$\mathcal{N}^{D}_{Q}(\lambda)$ $\leq\frac{|Q|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{\ell(Q)\sqrt{\lambda}}\right)_{+}.$
-This together with the monotonicity of Dirichlet eigenvalues implies that
-$\frac{|Q|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{\ell(Q)\sqrt{\lambda}}\right)_{+}$ $\geq\mathcal{N}^{D}_{Q}(\lambda)\geq\mathcal{N}^{D}_{Q\setminus\tilde{Q}}(\lambda)+\mathcal{N}^{D}_{\tilde{Q}}(\lambda)$
+Acknowledgments
+We wish to thank Prof.
+Rupert Frank and Prof.
+Simon Larson for bringing references [9, 12, 28] to us and several useful comments, and thank Dr.
+Xiang He for bringing the reference [32].
+R.
+Jiang wishes to thank Prof.
+Xiao Zhong for explaining to him Pólya’s conjecture and helpful discussions at University of Jyväskylä around 2010.
+R.
+Jiang was partially supported by NNSF of China (12526205 & 12471094), F.H.
+Lin was in part supported by the National Science Foundation Grant DMS2247773 and DMS1955249.